### v1

1 Introduction
High-precision predictions for particle colliders rely more than ever on the computation of multi-loop Feynman integrals. For this reason, the study of Feynman integrals and their analytic structure has become a major area of research in theoretical particle physics (see, e.g., [1, 2] for reviews). A central challenge in these computations is the reliable prediction of Feynman integral singularities as functions of kinematic parameters, without the need for a direct evaluation of the integral. In this paper, we tackle this problem by introducing [Uncaptioned image] (Singularities of Feynman Integrals Automatized), a Mathematica package designed for this purpose, which we release for public use:
SOFIA GitHub repository [3].
Differential equations.
The most common application is in computing Feynman integrals using the method of differential equations [4]. In the simplest cases, these equations take the form
$\mathrm{d}\vec{I}=\sum_{i}\mathbf{\Omega}_{i}\,\mathrm{d}\log(W_{i})\,\vec{I}\,,$ (1.1)
where $\vec{I}$ is a vector of basis Feynman integrals with $\mathrm{d}$ refers to the differential in the kinematic variables. On the right-hand side, $\mathbf{\Omega}_{i}$ are matrices with entries rational in the kinematics and the spacetime dimension $\mathrm{D}$. In fact, one often looks for bases in which all $\mathbf{\Omega}_{i}$ are linear in the dimensional-regularization parameter $\varepsilon=\frac{4-\mathrm{D}}{2}$ [5]. In this case, the resulting Feynman integrals $\vec{I}$ are often expressible in terms of multiple polylogarithms. The arguments of the logarithms, $W_{i}$, are known as the letters. Being able to predict these letters is currently one of the main bottlenecks in computations of polylogarithmic integrals: once $W_{i}$’s are known, one can fit the coefficients $\mathbf{\Omega}_{i}$ and consequently compute $\vec{I}$ using the differential equation together with some boundary conditions.
Many caveats and extra complications can arise in this story, but the motivation remains the same: We want to determine the set of all possible $W_{i}$’s known as the alphabet. This application is the main motivation for our work.
(a) 135 singularities (10m)
369 letters (84m)
SolverBound -> 1000
(b) 125 singularities (11m)
237 letters (70m)
SolverBound -> 100
(c) 275 singularities (10m)
? letters
SolverBound -> 500
Figure 1: Examples of cutting-edge diagrams and with timings (in minutes) to obtained singularities and letters. See [3] for expressions. Dashed lines denote massless particles and otherwise mass labels are displayed along the edges. The choice of SolverBound is determined by the maximum value for which the letters could be subsequently obtained. For diagram (c), although the singularity analysis finished in a reasonable time, the extraction of letters did not complete (hence the “?”) for any reasonable choice of SolverBound. For diagram (a), we used a machine with two Intel Xeon Gold CPUs and 512GB memory, while for (b) and (c) we used a MacBook Pro with an M3 chip and 16GB memory.
Local behavior near singularities.
Since problems involving analytic structure are notoriously subtle, let us be more precise about the type of questions that we aim to address. A singularity refers to a pole or a branch point in the (complexified) kinematic space. Let us collectively denote the kinematic variables as $\vec{s}\in\mathbb{C}^{\mathfrak{s}}$: they consist of the $\mathfrak{s}$ Mandelstam invariants and masses involved in the problem. While non-perturbatively scattering amplitudes might have a much richer singularity structure, in perturbation theory it is expected that singularities of the type
$I_{a}\sim f_{j}^{A}\,\log^{B}f_{j}\,,$ (1.2)
where $f_{j}=f_{j}(\vec{s})$ is a polynomial and $A$, $B$ are rational numbers. For example, $\sim\sqrt{f_{j}}\log f_{j}$ is one possibility. Note that this refers to a local behavior of the amplitude near the locus $f_{j}=0$. It is much simpler than the global behavior which has to be compatible with (1.2) for a list of all possible $f_{j}$’s. Indeed, it is known that Feynman integrals are expressed in terms of complicated functions such as multi-polylogarithms, elliptic functions, periods of higher-genus Riemann surfaces and Calabi–Yau manifolds, etc. [6, 7] even though their local behavior (1.2) remains simple. The goal of singularity analysis is to exploit this simplicity.
The connection between the local behavior near singularities (1.2) and the global structure of differential equations (1.1) is very simple. It must be that zeros, poles, and singularities of $W_{i}$’s are contained in the list of singularities $f_{j}$. Indeed, our goal will be to first compute a (possibly overcomplete) list of $f_{j}$’s. Then, we will reconstruct a list of possible $W_{i}$’s compatible with the aforementioned constraints.
Physical interpretation.
In the literature, the aforementioned singularities $f_{j}=0$ are also interchangeably known as Landau singularities or anomalous thresholds [8, 9, 10, 11]. The equation $f_{j}=0$ is a condition for “billiard ball scattering”: a network of classical on-shell particles following their classical, though possibly complexified, paths. The simplest case corresponds to positive energies and real angles [12], i.e., the physical region. But the interpretation extends to the rest of the physical sheet, and in fact any sheet, after relaxing positivity and reality conditions; see, e.g., [13].
Recent progress.
There has been an explosion of recent activity in the topic of determining computing singularities of Feynman integrals. Let us mention a couple of algorithms that have been implemented as parts of public packages and can handle state-of-the-art Feynman integrals. HyperInt implements a reduction algorithm that finds a superset of Landau singularities [14]. PLD.jl implements an algorithm based on toric geometry that finds a computable subset of Landau singularities [15, 16]. Both of these approaches work in the Schwinger-like parametrizations of Feynman integrals. A database comparing the two codes can be found in [17]. The package Baikovletter attempts to find symbol alphabets using Baikov representations [18]. We refer the readers to Sec. 2.2 for more detailed comparison and the introduction of [16] for a more complete list of references.
Importantly, it was recently observed that applying singularity analysis to the elastic unitarity constraint (relating the imaginary part of a diagram to its “cut” sub-diagrams) yields a powerful recursion formula for the calculation of Landau singularities in the loop momentum space [19]. This method enables efficient computation of the leading Landau singularities in two-particle reducible diagrams, successfully reproducing known results from the database [17] while also making new predictions, including all-loop examples.
In this paper, we extend this approach to any Feynman diagram. Building on the aforementioned advanced computational techniques and additional theoretical understanding, we propose an algorithm that offers a systematic way to study the full set of Landau singularities in general Feynman diagrams.
New contributions.
The main goal of [Uncaptioned image] is to streamline the analysis of singularities with minimal user input. The package comes with a companion Mathematica notebook containing 35 examples, where known results are reproduced and new predictions are given. This paper serves as a manual for [Uncaptioned image] .
The first main novelty comes from a new variant of Baikov representations [20, 21] adapted to minimize the number of integration variables and polynomials involved in the singularity analysis. [Uncaptioned image] implements an optimization technique that results in a minimal integral representation (according to a metric explained in Sec. 2.1) without any user input besides the Feynman diagram under consideration. We believe it could come useful in other aspects of Feynman analysis that rely on Baikov representations, such as integration-by-parts reduction or analyzing Feynman geometries [22, 23, 24, 25, 26, 27, 28, 29, 30, 31]. We illustrate the latter application in Sec. 4.2. Likewise, we extend our analysis to eikonalized Feynman integrals, such as those in post-Minkowskian expansions with an example in Sec. 4.3.
The second key innovation is the integration of advanced polynomial system solving methods for singularity analysis into the optimized Baikov representation. In particular, we implement a variant of the Fubini reduction algorithm [32] that results in an upper bound on the singularity locus. We also implement a Mathematica interface to PLD.jl, which allows one to seamlessly perform a numerical and symbolic analysis on a given diagram. We emphasize that this algorithm can be applied to any Euler integral and give a couple of examples of applications to cosmological and energy correlators in Sec. 4.4.
The third and final significant novelty is the implementation of graph-theoretical tools to break down and speed up the computation of singularities. It has long been observed that the symmetries of a Feynman diagram lead to an identification between Schwinger parameters of the leading physical sheet singularity [33, 34] (see [35] for a modern application in the context of the S-matrix bootstrap). In this work, we extend this idea by relating different sets of subleading singularities through the construction and identification of subtopologies that are connected via symmetry transformations accounting for variations in their kinematics. This procedure effectively mitigates the factorial growth in the number of subtopologies that would otherwise need to be analyzed.
Finally, we use the recent implementation of the publicly available packge Effortless [36] to provide an end-to-end function that takes a diagram as the input and provides a candidate symbol alphabet as the output. By “candidate” we mean that the alphabet might be overcomplete. An example is provided in Sec. 4.1.
We believe that [Uncaptioned image] establishes substantial new milestones in the singularity analysis of Feynman integrals, producing results that were previously unattainable. The ancillary files [3] contain several new predictions for singularities and alphabets, including those presented in Fig. 1, which exemplify how [Uncaptioned image] can be applied to next-frontier examples, namely, $HZ+2$ jets, $H+$jet, and multi-jets production through light- (for (a)) or heavy-quark (for (b) and (c)) loop-induced gluon fusion at the LHC.
We hope that [Uncaptioned image] serves as a tool to explore new avenues and sheds light on the analytic structures of perturbative scattering amplitudes that were previously inaccessible.
Outline.
We provide theoretical background in Sec. 2, including the discussion of different variants of the Baikov representation, singularity analysis of integrals, and converting from singularities to symbol alphabets. The manual of [Uncaptioned image] is given in Sec. 3 with an outline of all the functions and options. Sec. 4 focuses on examples and various applications. Finally, future directions are discussed in Sec. 5.
2 Theoretical background
The purpose of this section is to introduce three concepts: the Baikov representations, singularity analysis, and symbol alphabets.
2.1 Baikov representations
[Uncaptioned image] is based on the Baikov representation of Feynman integrals. In this section, we illustrate two variations of the Baikov representation: global and loop-by-loop. We will illustrate them directly on the simple example of the sunrise diagram. More complicated examples will be studied in Sec. 4, and a detailed derivation of the Baikov representation can be found in App. A. All the (loop-by-loop) Baikov representations given below can be generated with the [Uncaptioned image] package described in Sec. 3.3.
Normalization.
Throughout this paper and in the package, we use the following conventions for scalar Feynman integrals
$I_{\vec{\nu}}(\vec{s})=\frac{1}{(i\pi^{\mathrm{D}/2})^{L}}\int\frac{\mathrm{d}^{\mathrm{D}}\ell_{1}\,\mathrm{d}^{\mathrm{D}}\ell_{2}\cdots\mathrm{d}^{\mathrm{D}}\ell_{L}}{(q_{1}^{2}-m_{1}^{2})^{\nu_{1}}(q_{2}^{2}-m_{2}^{2})^{\nu_{2}}\cdots(q_{N}^{2}-m_{N}^{2})^{\nu_{N}}}\,,$ (2.1)
where $\mathrm{D}$ is the spacetime dimension, $\ell_{a}$ denote the loop momenta, and $q_{e}$ and $m_{e}$ are the momenta and masses of the internal edges. The $\nu_{i}$’s are the (integer) powers of the propagators. The list $\vec{s}$ collectively denotes the external kinematic variables such as Mandelstam invariants and masses.
General strategy.
The idea of Baikov representation is to translate the integration over the loop momenta $\ell_{a}$ into that over Lorentz products of the form $\ell_{a}\cdot p_{i}$ and $\ell_{a}\cdot\ell_{b}$, where $p_{i}$ denote the external momenta in all-incoming conventions. Schematically, we want to convert the integration measure
$\mathrm{d}^{\mathrm{D}}\ell_{1}\,\mathrm{d}^{\mathrm{D}}\ell_{2}\cdots\mathrm{d}^{\mathrm{D}}\ell_{L}\quad\mapsto\quad\mathcal{J}\,\mathrm{d}(\ell_{1}\cdot p_{1})\,\mathrm{d}(\ell_{1}\cdot p_{2})\cdots\mathrm{d}(\ell_{1}\cdot\ell_{1})\,\mathrm{d}(\ell_{1}\cdot\ell_{2})\cdots\,.$ (2.2)
The change of variables comes with a Jacobian which we called $\mathcal{J}$. We can go one step further after we notice that the new variables are related to inverse propagators $x_{i}=(\ell_{a}+\ldots+p_{i}+\ldots)^{2}-m_{i}^{2}$, which are always quadratic in the loop momenta, by a linear transformation. Therefore, we can express
$\mathrm{d}^{\mathrm{D}}\ell_{1}\,\mathrm{d}^{\mathrm{D}}\ell_{2}\cdots\mathrm{d}^{\mathrm{D}}\ell_{L}\quad\mapsto\quad\tilde{\mathcal{J}}\,\underbracket{\mathrm{d}x_{1}\,\mathrm{d}x_{2}\,\ldots\,\mathrm{d}x_{E}}_{\text{inverse propagators}}\,\underbracket{\mathrm{d}x_{E+1}\,\ldots\mathrm{d}x_{\mathcal{N}}}_{\text{ISPs}}\,,$ (2.3)
with yet another Jacobian $\tilde{\mathcal{J}}$. The caveat is that the number of variables $\mathcal{N}$ is in general larger than the number of propagators $E$. Therefore, $x_{1},\ldots x_{E}$ label inverse propagators, while $x_{E+1},\ldots,x_{\mathcal{N}}$ are the so-called irreducible scalar products (ISPs).
Global vs. loop-by-loop approaches.
There are a few ways to implement the change of variables (2.3). The general idea dates back to the S-matrix theory in the 1960s [37, 38, 39], although in the more recent literature on Feynman integrals parametrizations of this type are referred to as Baikov representations [20]. In the global Baikov representation, one transforms all variables at the same time. This results in the representation of the form
$I(\vec{s})=c\,\tilde{\mathcal{G}}(\vec{s})^{\tilde{\gamma}}\int_{\Gamma^{\prime}}\mathcal{G}(\vec{x};\vec{s})^{\gamma}\frac{\mathrm{d}^{N^{\prime}}\vec{x}}{x_{1}^{\nu_{1}}x_{2}^{\nu_{2}}\cdots x_{E}^{\nu_{E}}}\,,$ (2.4)
where $\vec{x}=(x_{1},\ldots,x_{E},\ldots,x_{N^{\prime}})$ are the integration variables with
$\mathcal{N}\mapsto N^{\prime}=\frac{L(L+1)}{2}+L\mathcal{E}\,,$ (2.5)
where $\mathcal{E}$ is the number of linearly-independent external momenta and $\vec{s}$ collectively denotes the external kinematic variables such as Mandelstam invariants and masses. The exponents $\gamma$, $\tilde{\gamma}$, as well as the prefactor $c$ are expressed in terms of the topology of the diagram (number of edges, etc). as well as the spacetime dimension $\mathrm{D}$. The derivation is given in App. A and the explicit expressions for all the ingredients in (2.4) are spelled out in (A.24).
The central object in the global Baikov representation is $\mathcal{G}(\vec{x},\vec{s})$ which arises as the general Jacobian for the change of variables. It can be expressed as a Gram determinant of a certain matrix. The integration contour $\Gamma^{\prime}$ is determined by $\vec{x}$ such that $\mathcal{G}(\vec{x},\vec{s})>0$. Likewise, $\tilde{\mathcal{G}}(\vec{s})$ is another Gram determinant that depends only on external kinematics and hence acts as an overall normalization. The derivation in App. A makes it clear that one can include different powers of propagators as well as non-trivial numerators with minimal modifications. For completeness, we review these aspects in App. A (see also [40, 41]).
The alternatives proposed by Frellesvig and Papadopoulos are loop-by-loop Baikov representations [22]. The idea is simply to Baikov parametrize each loop of the diagram sequentially in a given order. For each loop, we look for a shift in the loop momentum $\ell_{a}$ so that fewer invariants—such as $\ell_{a}\cdot p_{i}$ and $\ell_{a}\cdot\ell_{b}$—appear before changing variables from the $\ell$’s to the $x$’s. This procedure has the effect of decreasing the total number of integration variables $N\leqslant N^{\prime}$. Typically, the minimal/optimal $N$, is [42]
$\mathcal{N}\mapsto N=L+\sum_{a=1}^{L}\mathcal{E}_{a}\,,$ (2.6)
where $\mathcal{E}_{a}$ denotes the number of linearly independent external momenta associated with loop $\ell_{a}$. Note that in (2.6), $N$ scales linearly with $L$, whereas in (2.5), $N^{\prime}$ exhibits a quadratic dependence.
Loop-by-loop representations are not unique, as they depend on the exact change of variables. Therefore, one can distinguish between different variants, depending on how such choices are being made. Below, we will describe our optimization procedure. In any case, all the integral representations takes the general form:
$I(\vec{s})=c^{\prime}\int_{\Gamma^{\prime}}\tilde{\mathcal{G}}_{1}^{\tilde{\gamma}_{1}}\mathcal{G}_{1}^{\gamma_{1}}\tilde{\mathcal{G}}_{2}^{\tilde{\gamma}_{2}}\mathcal{G}_{2}^{\gamma_{2}}\cdots\tilde{\mathcal{G}}_{L}^{\tilde{\gamma}_{L}}\mathcal{G}_{L}^{\gamma_{L}}\frac{\mathrm{d}^{N}\vec{x}}{x_{1}^{\nu_{1}}x_{2}^{\nu_{2}}\cdots x_{E}^{\nu_{E}}}\,.$ (2.7)
Structurally, the only difference is that it features $L$ pairs of Gram determinants $\tilde{\mathcal{G}}_{a}$, $\mathcal{G}_{a}$ for every loop $a=1,2,\ldots,L$, instead of just one. Note that the order in which we label the loop momenta $\ell_{1},\ell_{2},\ldots,\ell_{L}$ plays a role. The integration contour is given by the intersection of regions defined by positivity conditions on internal-external Gram determinant ratios: $\Gamma^{\prime}=\bigcap_{l=1}^{L}\{\mathcal{G}_{l}/\tilde{\mathcal{G}}_{l}>0\}$. Details of the derivation are given in App. A.
Generalized cuts.
Due to complicated integration contours, Baikov representations are typically not used in the explicit evaluation of Feynman integrals. However, they are extremely useful for integral reduction, integration-by-parts identities, singularity analysis, and other applications that rely on the simplicity of the integrand. The main advantage comes from the application of generalized unitarity. Generalized cuts can be taken by simply replacing the integration contour with residues around the relevant propagators (similarly, unitarity or Cutkosky cuts can be taken after an extra modification of the integration contour). For example, the maximal cut is obtained by
$\int\mathrm{d}^{\mathcal{N}}\vec{x}\quad\mapsto\quad\frac{1}{(2\pi i)^{E}}\oint_{x_{1}=0}\oint_{x_{2}=0}\cdots\oint_{x_{E}=0}\int\mathrm{d}^{\mathcal{N}-E}\vec{x}\,,$ (2.8)
in either global or loop-by-loop Baikov approach (i.e., $\mathcal{N}=N\,,N^{\prime}$). This replacement exists if the original integration domain, $\Gamma$ or $\Gamma^{\prime}$, restricted to $x_{1}=x_{2}=\ldots=x_{\mathcal{N}}=0$ is non-empty (more on that is discussed in Sec. 3.3). Note that after taking the cut, $\mathcal{N}-E$ variables still have to be integrated over.
Implementation.
As mentioned above, Baikov representations are not unique. They depend on the order and labeling of the loop momenta $\ell_{a}$, as well as additional shifts.
We implement a search algorithm that optimizes for the most efficient loop-by-loop Baikov representation without any user input. It proceeds in two steps. First, when possible, the loops are ordered automatically so that the number of external momenta $\mathcal{E}_{a}$ (associated with each loop momentum $\ell_{a}$, for $1\leqslant a\leqslant L$) increases from the smallest to the largest. [^1] This strategy is inspired by the recursive elastic unitarity approach of [19], where the selection and ordering of cuts play a crucial role in determining the complexity of the singularity analysis. In [19], this selection was performed manually, whereas in the present work, the analogous loop-ordering process is fully automated. For example, for the diagram in Fig. 1(a), the algorithm selects the box (with $\mathcal{E}_{1}=3$) as the first loop and the hexagon (with $\mathcal{E}_{2}=5$) as the second. Second, the translation invariance of the integration measure is exploited by shifting the loop momenta, i.e., $\ell_{a}\mapsto\ell_{a}^{\prime}=\ell_{a}+\lambda_{a}$, to eliminate as many scalar products as possible. An explicit example of such shifts is provided in the sunrise example below.
Comparison with BaikovPackage.
Recently, a semi-automatized Mathematica package BaikovPackage for generating standard and loop-by-loop Baikov representations of Feynman integrals was released [21]. Unlike [Uncaptioned image] , which by default selects the loop-by-loop representation minimizing the number of integration variables $N$, BaikovPackage requires the user to make that choice and specify the order and labeling of the loop momenta $\ell_{a}$. Consequently, the outputs may differ depending on these selections.
Example.
Let us illustrate the differences between the global and loop-by-loop approaches on (arguably) the simplest example: the sunrise diagram. To make expressions shorter, we take all internal masses equal $m$ and external momentum $p^{2}=s$. In this case, we have only three propagators, $E=3$, with:
$x_{1}=\ell_{1}^{2}-m^{2}\,,\qquad x_{2}=\ell_{2}^{2}-m^{2}\,,\qquad x_{3}=(p-\ell_{1}-\ell_{2})^{2}-m^{2}\,.$ (2.9)
As predicted by (2.5) with $L=2$ and $\mathcal{E}=1$, expanding these propagators out, the result depends on $N^{\prime}=5$ internal kinematic invariants: $\ell_{1}\cdot p$, $\ell_{2}\cdot p$, $\ell_{1}\cdot\ell_{2}$, $\ell_{1}^{2}$, $\ell_{2}^{2}$, in addition to the external parameters $\vec{s}=(s,m^{2})$. The $5$ internal variables can be traded for $\vec{x}=(x_{1},x_{2},x_{3},x_{4},x_{5})$ by a linear change of variables, where $x_{4}$ and $x_{5}$ are arbitrarily-chosen ISPs, say
$x_{4}=\ell_{1}\cdot p\,,\qquad x_{5}=\ell_{2}\cdot p\,.$ (2.10)
This procedure results in the global Baikov parametrization (2.4) with
$\tilde{\mathcal{G}}(s,m^{2})=s\,,\quad\mathcal{G}(\vec{x},s,m^{2})=\det\left[\begin{array}[]{ccc}m^{2}{+}x_{1}&*&*\\ \frac{1}{2}\left(x_{3}{-}x_{1}{-}x_{2}{-}m^{2}{-}s\right){+}x_{4}{+}x_{5}&m^{2}{+}x_{2}&*\\ x_{4}&x_{5}&s\\ \end{array}\right]\,,$ (2.11)
where the “*” denote symmetric entries.
In contrast, loop-by-loop representations first parametrize all invariants associated with the first loop momentum $\ell_{1}$. They are $\ell_{1}\cdot(\ell_{2}-p)$ and $\ell_{1}^{2}$. Notice that the former is counted as one variable: $\ell_{1}\cdot\ell_{2}$ and $\ell_{1}\cdot p$ appear in the expression for the Feynman integral only in the combination $\ell_{1}\cdot(\ell_{2}-p)$. Hence, integrating out the first loop momentum would give a function of $(\ell_{2}-p)^{2}$, which is treated as external kinematics from this point of view. When it comes to parameterizing the second loop momentum $\ell_{2}$, we are thus left with only two kinematic invariants: $(\ell_{2}-p)^{2}$ and $\ell_{2}^{2}$. Consequently, the total number of variables required is $N=4$. This is consistent with (2.6) for $L=2$ loops and $\mathcal{E}_{1}=\mathcal{E}_{2}=1$. Notably, this approach reduces the total number of variables compared to the global Baikov parametrization, which requires $N^{\prime}=5$ variables.
Equivalently, we could have first relabeled $\ell^{\prime}_{2}=\ell_{2}-p$, which turns the inverse propagators into
$x_{1}=\ell_{1}^{2}-m^{2}\,,\qquad x_{2}=(\ell_{2}^{\prime}+p)^{2}-m^{2}\,,\qquad x_{3}=(\ell_{1}+\ell_{2}^{\prime})^{2}-m^{2}\,.$ (2.12)
This makes it clear that only four invariants are needed: $\ell_{1}$, $\ell_{2}^{\prime 2}$, $\ell_{1}\cdot\ell_{2}^{\prime}$, and $p\cdot\ell_{2}^{\prime}$. Taking $x_{4}=\ell_{2}\cdot p$ as the only ISP, the resulting representation has
$\tilde{\mathcal{G}}_{1}$ $=m^{2}+s+x_{2}-2x_{4},$ $\mathcal{G}_{1}$ $=\det\left[\begin{array}[]{cc}m^{2}+x_{1}&*\\ \frac{x_{3}-x_{1}-x_{2}-s-m^{2}}{2}+x_{4}&m^{2}+s+x_{2}-2x_{4}\end{array}\right],$ (2.15)
$\tilde{\mathcal{G}}_{2}$ $=s,$ $\mathcal{G}_{2}$ $=\det\left[\begin{array}[]{cc}m^{2}+x_{2}&*\\ x_{4}&s\end{array}\right].$ (2.18)
We now proceed to explain how the singularity analysis is carried out in Baikov space.
2.2 Singularity analysis: Fubini reduction algorithm
As a mathematical problem, singularity analysis amounts to analyzing the topology of the union of hypersurfaces
$\mathcal{U}(\vec{x},\vec{s})\equiv\{g_{1}(\vec{x};\vec{s})=0\}\cup\{g_{2}(\vec{x};\vec{s})=0\}\cup\cdots\cup\{g_{k}(\vec{x};\vec{s})=0\}\,,$ (2.19)
in the complex space $\mathbb{C}^{N}$. Here, $\vec{x}\in\mathbb{C}^{N}$ denotes the integration variables and $\vec{s}$ denotes the collection of all the parameters, such as the kinematic invariants and the masses. Each of the $k$ polynomials $g_{i}(\vec{x};\vec{s})$ describes a singularity of the integrand (such as a pole) or an integration boundary. A singularity can arise only if the topology of (2.19) changes, e.g., when a mass shell degenerates into a lightcone or two mass shells pinch.
Example.
For instance, loop-by-loop Baikov representations have
$(g_{1},g_{2},\ldots,g_{2L-1},g_{2L},g_{2L+1},\ldots,g_{2L+E})=(\mathcal{G}_{1},\tilde{\mathcal{G}}_{1},\ldots,\mathcal{G}_{L},\tilde{\mathcal{G}}_{L},x_{1},\ldots,x_{E})\,.$ (2.20)
The first $2L$ polynomials are the internal and external Gram determinants, while the last $E$ are the inverse propagators (and possibly ISPs). We do not need any extra polynomial for the integration boundaries, since they are already described in terms of the Gram determinants.
As another example, on the maximal cut, we need $E$ fewer polynomials:
$(g_{1},g_{2},\ldots,g_{2L-1},g_{2L})=(\mathcal{G}_{1},\tilde{\mathcal{G}}_{1},\ldots,\mathcal{G}_{L},\tilde{\mathcal{G}}_{L})\Big{|}_{x_{1}=x_{2}=\ldots=x_{E}=0}\,.$ (2.21)
In either case, one can add $x_{E+1},\ldots,x_{N}$ to the list if one allows for negative powers of ISPs.
Singularity analysis.
A conventional way to phrase this problem is to compute the Jacobian of the union of hypersurfaces:
$J=\left[\nabla g_{1}\;\nabla g_{2}\;\ldots\;\nabla g_{k}\right]\,,$ (2.22)
where $\nabla=\left[\partial_{x_{1}}\;\partial_{x_{2}}\;\ldots\;\partial_{x_{N}}\right]^{\intercal}$, and ask for what values of $\vec{s}$ the rank of $J$ changes; see, e.g., [43, Ch. 9]. In other words, we are asking for at least one null vector $\vec{\beta}=\left[\beta_{1},\beta_{2},\ldots,\beta_{k}\right]$ of $J$. (When applied to the loop-momentum representation (2.1) with $g_{i}=q_{i}^{2}-m_{i}^{2}$, $\beta$’s have the physical interpretation of Schwinger proper times). This requirement can be written as the following system of equations:
$\beta_{1}\nabla g_{1}(\vec{x},\vec{s})+\beta_{2}\nabla g_{2}(\vec{x},\vec{s})+\ldots+\beta_{k}\nabla g_{k}(\vec{x},\vec{s})=0\,,$ (2.23)
where $\vec{x}\in\mathcal{U}$. We call the codimension-1 part of the solution set $\mathcal{S}(\vec{s})$ and write it as
$\mathcal{S}(\vec{s})=\{f_{1}(\vec{s})=0\}\cup\{f_{2}(\vec{s})=0\}\cup\cdots\cup\{f_{K}(\vec{s})=0\}\,.$ (2.24)
The goal of singularity analysis is to convert $\mathcal{U}$ into $\mathcal{S}$. We have essentially converted a topological problem into an algebraic one. It has a certain combinatorial aspect since we can classify the solutions based on the number of $\{g_{i}=0\}$ surfaces that intersect at $\vec{x}$. The problem becomes even more difficult when $\mathcal{U}$ has a complicated topology since one has to take into account its non-normal crossings through blow-ups.
Singularity analysis of this kind applied to Feynman integrals dates back to the pioneering work of Bjorken [8], Landau [9], and Nakanishi [10] who wrote systems of polynomial equations (2.23) in various parametrizations. These are nowadays known as Landau equations and serve as a prototype for the analysis of singularities. This analysis was made rigorous through the work of Pham, Thom, Whitney, and others with the application of algebraic geometry tools to finite Feynman integrals; see [44] for a review. While mathematically rigorous, these methods tend to explode in computational complexity in anything but the simplest examples; see, e.g., [45] for a recent implementation in Macaulay2. Hence, recent resurgence of interest in this topic aims to make the singularity locus more computable.
Numerical approaches.
Let us briefly outline the scope of numerical tools that are currently available for cutting-edge Feynman integrals. The Julia package PLD.jl [15, 16] uses techniques from toric and polyhedral geometry to give a computable subset of the singularity locus $\mathcal{S}$. It also implements the Euler characteristic test, which is another way to probe the topology of the space $\mathcal{U}$. If for a given set of parameters $\vec{s}_{\ast}$, the absolute value of the Euler characteristic of
$(\mathbb{C}^{\ast})^{N}\setminus\mathcal{U}(\vec{x},\vec{s})\,,$ (2.25)
drops compared to its generic value, then $\vec{s}_{\ast}$ belongs to the singular locus. This test provides a quick way to determine whether a given $\vec{s}_{\ast}$ belongs to the singular locus $\mathcal{S}$. [Uncaptioned image] provides an interface to both functionalities of PLD.jl.
However, the package PLD.jl is based on toric geometry and hence uses =$\mathbb{C}^{\ast}=\mathbb{C}\setminus\{0\}$ instead of $\mathbb{C}$. This is appropriate for parametric representations of Feynman integrals, but in the Baikov representation it has the effect of introducing all $\{x_{i}=0\}$ in the hypersurface union $\mathcal{U}$. This is equivalent to putting ISPs with negative exponents.
Analytical approaches.
Let us now outline the available analytical approaches that lie at the heart of [Uncaptioned image] . In contrast to the aforementioned numerical approach, the goal will be to analytically compute a superset of the singularity locus $\mathcal{S}$. In other words, we expect to encounter certain “fictitious” polynomials $f_{i}(\vec{s})$ that do not satisfy the original system of equations (2.23). However, as we explain below, our algorithm is designed to minimize their occurrence.
Let us describe the Fubini reduction algorithm, which is a variation of [46, Sec. 4.2]. It circumvents the need to solve systems of equations such as (2.23). Instead, it proceeds by eliminating one variable $x_{i}$ at a time. We thus construct the sets of polynomials $\mathcal{L}$, $\mathcal{L}_{x_{i}}$, $\mathcal{L}_{x_{i},x_{j}}$, etc. recursively, where the subscript refers to the variables we already eliminated. At each step, we define
$[\mathcal{L}_{x_{i},\ldots,x_{j}}]_{x_{k}}\equiv\bigcup_{g\,\in\,\mathcal{L}_{x_{i},\ldots,x_{j}}}\!\!\!\!\{D_{x_{k}}(g),\,R_{x_{k}}(\infty,g)\}\;\cup\bigcup_{g,\hat{g}\,\in\,\mathcal{L}_{x_{i},\ldots,x_{j}}}\!\!\!\!\{R_{x_{k}}(g,\hat{g})\}\,.$ (2.26)
On the right-hand side, $D_{x_{k}}(g)$ refers to the discriminant of $f$ with respect to $x_{k}$, and $R_{x_{k}}(g,\hat{g})$ is the resultant of $g$ and $\hat{g}$ with respect to $x_{k}$. Writing the polynomial $g=c_{0}+c_{1}x_{k}+\ldots+c_{d}x_{k}^{d}$, the resultant $R_{x_{k}}(\infty,g)=c_{d}$ equals to the coefficient of the highest power. The set $[\mathcal{L}_{x_{i},\ldots,x_{j}}]_{x_{k}}$ might consist of reducible polynomials or repeated entries. We hence define $\mathcal{L}_{x_{i},\ldots,x_{j},x_{k}}=\text{irred}[\mathcal{L}_{x_{i},\ldots,x_{j}}]_{x_{k}}$ containing only the unique irreducible polynomials.
The three contributions in (2.26) have the following interpretations. $D_{x_{k}}(g)$ corresponds to self-pinching of $g=0$, $R_{x_{k}}(\infty,g)$ gives the pinch at infinity, while $R_{x_{k}}(g,\hat{g})$ is associated with a pinch between two curves $g=0$ and $\hat{g}=0$.
One strategy in computing the singularity locus would be therefore to start with
=$\mathcal{L}=\{g_{1},g_{2},\ldots,g_{k}\}\,.$ (2.27)
and then perform the reduction in some order:
$\mathcal{L}\mapsto\mathcal{L}_{x_{1}}\mapsto\mathcal{L}_{x_{1},x_{2}}\mapsto\ldots\mapsto\mathcal{L}_{x_{1},x_{2},\ldots,x_{N}}\,.$ (2.28)
The result $\mathcal{L}_{x_{1},x_{2},\ldots,x_{N}}$ consists of polynomials that depend on $\vec{s}$ only. In general, it is a (large) superset of the singularity locus $\mathcal{S}$. However, the result of the computation depends on the specific order of the variables we chose in the reduction.
We refer to the difference between $\mathcal{L}_{x_{1},x_{2},\ldots,x_{N}}$ and $\mathcal{S}$ as fictitious singularities. They are artifacts of the reduction procedure and have no physical meaning. In general, we are not concerned with catching a few fictitious singularities, since the goal is to find a superset of $\mathcal{S}$ anyway. However, it is beneficial to minimize their number.
The Fubini reduction algorithm instructs us to scan over all choices of $\mathcal{L}_{x_{1},x_{2},\ldots,x_{N}}$ and take the intersection of their results, that is
$\mathcal{L}_{\{x_{1},x_{2},\ldots,x_{N}\}}=\bigcap_{\sigma\in S_{N}}\mathcal{L}_{\sigma(x_{1},x_{2},\ldots,x_{N})}\,,$ (2.29)
where $S_{N}$ denotes the permutation set of $N$ labels (two polynomials are considered equivalent for the purposes of intersection if their vanishing locus is the same). In practice, a lot of the intermediate results can be recycled between each permutation. Hence, the optimal recursion is
$\mathcal{L}_{\{x_{1},\ldots,x_{k-1},x_{k}\}}=\bigcap_{i=1}^{k}\,\text{irred}[\mathcal{L}_{\{x_{1},\ldots,x_{k-1},x_{k}\}\setminus x_{i}}]_{x_{i}}\,,$ (2.30)
where each [$[\mathcal{L}_{\{x_{1},\ldots,x_{k-1},x_{k}\}\setminus x_{i}}]_{x_{i}}$ is defined as in (2.26). The central difference between (2.29) and (2.30) is that the former intersects $N!$ sets of proposed singularities at the very end, while the latter does so at the intermediate stages. This comes with enormous computational benefit and provides a slimmer superset of the singular locus $\mathcal{S}$, often coinciding with $\mathcal{S}$. The equation (2.30) is implemented in [Uncaptioned image] as the default solver option FastFubini.
Brown [32] and Panzer [47] have also proposed more sophisticated versions of the Fubini algorithm known as compatibility graph reduction. They are implemented in the Maple package HyperInt [14]; see [16, App. A] for a review. We found that our implementation of this approach was slower than (2.30) in practical applications, and is therefore not shipped with v1.0.0 of [Uncaptioned image] .
Other approaches.
Other approaches to determining symbol alphabets based on Landau analysis exist; for example, utilizing momentum twistor techniques or Schubert analysis; see, e.g., [48, 49, 50] for recent developments and more exhaustive lists of references. Out of the algorithms implemented as computer codes, the Mathematica package Baikovletter [51, 52, 18] uses Baikov representations to attempt constructing symbol letters. Instead of a reduction algorithm, it aims to return a list of candidate symbol letters directly. There are instances in which the resulting alphabet is not complete; see, e.g., [53, 54].
2.3 From singularities to letters
If a Feynman integral is sufficiently simple, it can admit a representation as a linear combination of multiple polylogarithms. In such cases, there is a great interest in knowing a priori the list of possible symbol letters $W_{i}$, which translate to arguments of these polylogarithms; see, e.g., [55, 56]. The collection of all independent letters is known as the symbol alphabet $\mathbb{A}$.
Before we move on, we note that even after identifying that a family of Feynman integrals includes integrals evaluating to non-polylogarithmic functions, knowledge of $\mathbb{A}$ remains valuable. This is because integrals in subsectors typically remain polylogarithmic (see [23, 57, 29] for explicit examples).
Even letters.
As reviewed in the introduction, letters are closely related to singularities. By definition, the zeros, poles, and singularities of letters are singularities of Feynman integrals. The simplest example is $\log(W_{i})$, which has branch points at $W_{i}=0$ and $W_{i}=\infty$. Hence the symbol alphabet must at least contain all the singularities themselves:
$(W_{1},W_{2},\ldots,W_{K})=(f_{1},f_{2},\ldots,f_{K})=\mathbb{A}^{\text{even}}\,.$ (2.31)
(One can also consider ratios of polynomials of the form $f_{j}/f_{k}$.) These are sometimes called even letters and certainly constitute a subset of the alphabet $\mathbb{A}^{\text{even}}\subseteq\mathbb{A}$. The question therefore becomes how to leverage this information to recover the remaining part of $\mathbb{A}$.
Odd letters.
We start by assuming the remaining letters, which we call odd, take the general form
$W_{i}=\frac{P_{i}+R_{i}\sqrt{Q_{i}}}{P_{i}-R_{i}\sqrt{Q_{i}}}\,,$ (2.32)
where $P_{i}=P_{i}(\vec{s})$, $Q_{i}=Q_{i}(\vec{s})$, and $R_{i}=R_{i}(\vec{s})$ are some homogeneous polynomials in $\vec{s}$, such that the mass dimensions match. We assume that $Q_{i}$ is square-free, so that it does not have any factors outside the square root. This is of course just an ansatz, and one can extend it to more complicated forms if necessary. We can now apply the same analysis as before. First of all, $W_{i}$ itself has a singularity at $Q_{i}=0$. Therefore, we must have
$Q_{i}=c_{i}\prod_{j=1}^{K}f_{j}^{e_{ij}}\,,$ (2.33)
where $c_{i}\in\mathbb{Q}$ is a rational number and $e_{ij}\in\mathbb{Z}_{\geqslant 0}$ are non-negative integer powers. Next, we need to check that zero and poles are singularities, which leads to a similar constraint:
$(P_{i}+R_{i}\sqrt{Q_{i}})(P_{i}-R_{i}\sqrt{Q_{i}})=P_{i}^{2}-R_{i}^{2}Q_{i}=c^{\prime}_{i}\prod_{j=1}^{K}f_{j}^{e^{\prime}_{ij}}\,,$ (2.34)
for some new sets of constants $c_{i}^{\prime}\in\mathbb{Q}$ and exponents $e_{ij}^{\prime}\in\mathbb{Z}_{\geqslant 0}$.
Differential equation.
As another perspective on the same problem, let us insert the odd letter ansatz (2.32) into the differential equation (1.1). The relevant term is
$\mathrm{d}\log(W_{i})=\frac{P_{i}R_{i}\mathrm{d}Q_{i}-2(\mathrm{d}P_{i}R_{i}-P_{i}\mathrm{d}R_{i})Q_{i}}{(P_{i}^{2}-R_{i}^{2}Q_{i})\sqrt{Q_{i}}}\,.$ (2.35)
Hence, indeed, the only singular points of the differential equation are located at the positions of Landau singularities $f_{j}=0$ for all possible $j$.
Therefore, knowing the symbol alphabet is not essential for constructing an ansatz for the differential equation (1.1) based solely on the Landau singularities $f_{j}$. However, such approach would lead to an unwieldy polynomial in the numerator. The symbol alphabet provides a practical way to reduce this ansatz into a more manageable form.
Effortless implementation.
There are a few different strategies to finding odd letters. To our knowledge, they were first explored in [58]. We will follow [59, Sec. 3.4], since this is the algorithm implemented in the publicly available package Effortless [36, 60]. The package assumes that a list of all candidate polynomials $Q_{i}$, $R_{i}$ in (2.32) and constants $c_{i}^{\prime}$ from (2.34) are given. It then attempts to construct all polynomials $R_{i}^{2}Q_{i}+c_{i}^{\prime}\prod_{j=1}^{K}f_{j}^{e_{ij}^{\prime}}$ that are a perfect square $P_{i}^{2}$. This can be done by scanning over all choices of $e_{ij}^{\prime}$. There is only a finite number of choices, since the set of singularities $f_{j}$ is finite and the exponents are limited because the whole expression has to have a uniform mass dimension. Repeating the same exercise for all choices of $Q_{i}$, $R_{i}$, and $c_{i}^{\prime}$ results in the proposed odd letters. Further checking for linear dependence, one obtains a candidate odd alphabet $\mathbb{A}^{\text{odd}}$ and hence also $\mathbb{A}=\mathbb{A}^{\text{even}}\cup\mathbb{A}^{\text{odd}}$.
In practice, we consider all $Q_{i}$ of the form
$\sqrt{Q_{i}}=\sqrt{g_{a}}\sqrt{h_{b}}\sqrt{f_{k}}\sqrt{f_{l}}\,,$ (2.36)
for every pair $k,l=1,2,\ldots,K$ with $R_{i}=1$ and $c_{i}^{\prime}=\pm 4$ (the default option in Effortless), as well as a possible user-specified set of polynomials $(g_{1},g_{2},\ldots,g_{A})$ and $(h_{1},h_{2},\ldots,h_{B})$ with $a=1,2,\ldots,A$ and $b=1,2,\ldots,B$, see Sec. 3.3.1 for details. Hence, in total, there are $A\times B\times K^{2}$ possibilities checked. The odd letter ansatz in [Uncaptioned image] v1.0.0 is formulated using the root structure defined in (2.36).
There is a handful of examples of Feynman integrals in which letters more complicated than the ansatz (2.32) exist, for example those with nested square roots [61, 62, 63]. These are currently not implemented in Effortless.
3 Manual
This section covers how to install [Uncaptioned image] and its dependencies, summarizes the key functions, and provides complete documentation.
3.1 Installation
As mentioned above, [Uncaptioned image] is hosted in a GitHub repository at [3], where a compressed file with the latest version can be downloaded or cloned. The package requires a working installation of Mathematica. It was developed and tested using the Mathematica versions 13.2–14.0 and can be loaded by running the command:
1Get["path/to/SOFIA.m"]
Alternatively, one may add the path to the package to the variable \$ Path, in which case the loading of the package amounts to calling Get["SOFIA.m"]. Running this command by itself uses the default options:
1SOFIAoptionFiniteFlow = False;
2SOFIAoptionJulia = False;
When an Internet connection is available, the Effortless package is automatically loaded from GitHub [60], ensuring that [Uncaptioned image] always uses the latest version. If needed, users can overwrite this path with SOFIAoptionEffortlessPath = "path/to/Effortless.m" to point to, say, a locally cloned copy. To speed up the alphabet reconstruction with Effortless, we recommend enabling FiniteFlow [64] by setting SOFIAoptionFiniteFlow = True before running Get["SOFIA.m"]. Similarly, the Julia dependencies can be activated by setting SOFIAoptionJulia = True, although this option is typically unnecessary (more on that below). A short guide for setting up an optional link between Mathematica and Julia can be found in App. B.
3.2 Summary for people with busy schedules
We illustrate how to use the main functionality of [Uncaptioned image] , which converts a diagram into a proposed list of singularities or candidate symbol letters depending on the options used.
Input.
The input diagram is specified through the list of its edges and nodes. The list edges consists of tuples { { i,j} ,m} specifying that a vertex i and j are connected through a propagator with mass m. The order does not matter. The nodes is a list of tuples { i,M} specifying that an external momentum $p_{i}$ with $p_{i}^{2}=\mathtt{M}^{2}$ is attached to the vertex i. The order does matter here since we call the external momenta $p_{1},p_{2},\ldots$ in the order of appearance.
As an example, the “parachute” diagram shown at the top of Fig. 2(a) corresponds to
1edges = {{{1, 2}, m}, {{1, 2}, m}, {{1, 3}, m}, {{2, 3}, m}};
2nodes = {{1, $\mathtt{M_{1}}$}, {2, $\mathtt{M_{2}}$}, {3, $\mathtt{M_{3}}$}};
The code automatically assigns independent Mandelstam invariants of the form sij… $=(p_{i}+p_{j}+\ldots)^{2}$ in the cyclic basis. For example, $6$-particle Mandelstam invariants are s12, s23, s34, s45, s56, s16, s123, s234, s345 (we work in $\mathrm{D}$ dimensions). Alternatively, one can run the command:
1{edges, nodes} = FeynmanDraw
which opens a window in which one can draw the diagram directly. The output of this command is the pair { edges, nodes} with generic masses. One can visually confirm that the diagram is correctly specified by plotting it with:
1FeynmanPlot[{edges, nodes}]
which reproduces the top of Fig. 2(a).
Note that the input masses in { edges, nodes} typically follow the format $\mathtt{M}$, $\mathtt{M}{}_{\mathtt{i}}$ (equivalent to Subscript[M, i]), $\mathtt{m}$, or $\mathtt{m}{}_{\mathtt{i}}$. However, as long as the Julia dependencies are not enabled, users can choose their own variable names. The only effect is that the mass variables will not be automatically homogenized in terms of mass squared in the output of the singularity analysis performed by the functions we now describe.
SOFIA command: Most important options and output format.
With this input, running [Uncaptioned image] amounts to:
1SOFIA[{edges, nodes}]
This command produces a list of candidate singularities. For the parachute diagram with edges and nodes given above, this yields the list at the top of Fig. 2(b). The detailed structure of the output—including variations produced under different options—will be discussed later in this section.
The function SOFIA[] comes with many options that can customize the computation depending on the specific application. The most important options (and their default values) are:
1SOFIA[{edges, nodes}, FindLetters -> False
2 , MaxCut -> True
3 , IncludeSubtopologies -> True
4 , Symmetries -> True
5 , SolverBound -> 100]
The complete list of options is given in Sec. 3.3.
Constructing candidate alphabets.
When the FindLetters -> True option is enabled, SOFIA[] first extracts the kinematic singularities and then uses the publicly available package Effortless [60] to generate all candidate letters, both even and odd. The structure of the output is then a nested list, with independent even and odd letters in the first and second entries, and square roots in the third entry, respectively:
1{{Even letters}, {Odd letters}, {Square roots}}
[…]
Maximal cuts and subtopologies.
The option MaxCut -> True, enabled by default, computes the leading singularities, which we take to be the singularities on the maximal cut (2.8). When IncludeSubtopologies -> True is also enabled, the calculation iterates over all subtopologies, therefore computing also subleading singularities.
[…]
Symmetries.
When IncludeSubtopologies -> True, the role of the option Symmetries -> True is to reduce the total number of subtopologies analyzed, by identifying those that are symmetric under a map between kinematic scales. These include transformations between internal and external masses, and Mandelstam invariants of the subtopologies. After the minimal set of subtopologies is analyzed, the singularities of the remaining ones are obtained via replacement rules. Thus, the option Symmetries -> True should reproduce the results as if all diagrams were considered, but with significantly improved efficiency. This is particularly valuable in high-multiplicity cases (i.e., six or more external particles), where it generally speeds up the SOFIA[] command by at least a factor of 2 to 10 in the examples considered. In general, its effect on reducing runtime is expected to increase with the number of external legs and loops. The specifics of how the symmetry map between subtopologies is constructed using the Symmetry[] function is discussed later in Sec. 3.3.4 below.
[…]
Bound in the number of monomials.
[…]
3.3 Documentation of functions
The purpose of this section is to provide a more comprehensive documentation of all the most important functions included in [Uncaptioned image] and their options. All these functions are internally memorized to improve performance, and in Sec. 4 we provide concrete examples that illustrate their usage.
3.3.1 SOFIA command
The singularities and associated alphabets are computed by the central function SOFIA[], whose high-level behavior and performance may depend on several user-specified options. Below, we list all its available options together with the default values:
1SOFIA[{edges, nodes}, FindLetters -> False
2 , IncludeSubtopologies -> True
3 , IncludeISPs -> False
4 , MaxCut -> True
5 , LoopEdges -> Automatic
6 , ShowPossiblyDegenerate -> False
7 , Symmetries -> True
8 , ShowHistoryGraph -> False
9 , UnclogTime -> 10^17
10 , StartAtSubtopology -> 1
11 , EndAtSubtopology -> -1
12 , Substitutions -> {}
13 , Solver -> FastFubini
14 , SolverBound -> 100
15 , FactorResult -> True
16 , PLDMethod -> sym
17 , PLDHomogeneous -> true
18 , PLDHighPrecision -> false
19 , PLDCodimStart -> -1
20 , PLDFaceStart -> 1
21 , PLDRunASingleFace -> false
22 , AddSing -> {}
23 , DoubleRoots -> {}
24 , AddLetters -> {}]
Lines 1–12 define options related to the Baikov representation, graph specifications, and scanning over subtopologies.
Baikov options.
[…]
ISPs.
[…]
Skipping and targeting subtopologies.
For examples that are still computationally expensive with IncludeSubtopologies -> True and Symmetries -> True, the user may want to set an upper bound on the time T (in seconds) taken by [Uncaptioned image] to perform the singularity analysis. This can be done using the options UnclogTime -> T. The default value is the age of the Universe. If the analysis of a (sub)topology exceeds time T, the code will record the position of the clogging components, allowing the user to inspect them individually later and optimize performance by adjusting other options.
[…]
Subtopology history graph.
When IncludeSubtopologies -> True, the option ShowHistoryGraph can be used to print a top-bottom graph (laid out based on its spanning tree) whose nodes represent (sub)topologies, with edges connecting those related by edge contractions. If ShowHistoryGraph -> True, both the diagrams and the corresponding singularities are displayed at each node (see Figs. 2(a) and 2(b), respectively); if ShowHistoryGraph -> SingOnly, only the singularities are shown.
Figure 2: Example output of SOFIA[{ edges, nodes} , ShowHistoryGraph -> True, SolverBound -> Infinity] applied to the parachute diagram with edges = { { { 1,2} ,m} , { { 1,2} ,m} , { { 2,3} ,m} , { { 1,3} ,m} } and nodes = { { $\mathtt{M_{1}}$} , { $\mathtt{M_{2}}$} , { $\mathtt{M_{3}}$} } . Panel (a) illustrates how diagrams are connected through edge contractions (with tadpoles removed), while panel (b) shows the corresponding candidate singularities. As explained in the main text, SOFIA[] uses the shorthands $\mathtt{mm=m^{2}}$ and $\mathtt{MMi=M_{i}^{2}}$ in its outputs. Note that the singularities in (b) are in one-to-one correspondence with those in the database [17]—that is, there are no fictitious singularities in this example.
Numerical substitutions.
[…]
Degenerate cuts.
[…]
Lines 13–21 give options that control the solving strategy.
Solver options.
Within [Uncaptioned image] , there are two available solvers: FastFubini (which implements the analytic procedure outlined in Sec. 2.2) and momentumPLD (which instead uses the numerical approach of Sec. 2.2). The former relies solely on Mathematica functions, while the latter requires the working Julia dependencies detailed in App. B. All the options PLDname relate specifically to momentumPLD (lines 13–18) are described under the same name as in [65]. Finally, the option FactorResult -> False can be used to skip the factorization into irreducible components of $\mathcal{S}(\vec{s})$ in (2.24), which, for large examples, can require both a long runtime and a large amount of memory.
[…]
Lines 22–24 set options for Effortless, which are relevant only when FindLetters -> True is enabled.
Effortless options.
At times, users might want to supplement the singularities computed by the SOFIA[] command with additional ones. For example, if singularities of subtopologies are already known or anticipated and IncludeSubtopologies -> False is used. To do so, one can use the option AddSing -> { p1,p2,...} to append the specified polynomials to the singularities found by SOFIA[].
By default (DoubleRoots -> { } ), SOFIA[] uses $g_{a}=h_{b}=1$ in (2.36) to construct the candidate odd alphabet $\mathbb{A}^{\text{odd}}$ from the ansatz in (2.32). The DoubleRoots option allows more flexibility. Specifying DoubleRoots -> { { fm,..., fn} ,{ fp,..., fq} } forces SOFIA[] to consider in (2.36) the polynomials provided by the user (yet typically belonging to (2.31)): { fm,..., fn} for $g_{a}$ and { fp,..., fq} for $h_{b}$. For instance, setting DoubleRoots -> { { f1, f2} ,{ f1,f3,f4} } ensures that $g_{a}=\mathtt{f1},\mathtt{f2}$ and $h_{b}=\mathtt{f1},\mathtt{f3},\mathtt{f4},\mathtt{f4}$ are considered in building the candidate $\mathbb{A}^{\text{odd}}$. Selecting DoubleRoots -> { { fm, ..., fn} , All} instructs SOFIA[] to assign the list { fm,..., fn} to $g_{a}$ while using all $f_{i}$ from (2.31) (which are internally constructed by SOFIA[]) for $h_{b}$.
In contrast, choosing DoubleRoots -> All directs SOFIA[] to apply all $f_{i}$ from (2.31) to both $g_{a}$ and $h_{b}$. In most cases, using the “sledgehammer” option DoubleRoots -> All comes at the cost of using significantly more computational resources. An example using the DoubleRoots option is discussed in Sec. 4.1.
[…]
3.3.2 SOFIADecomposeAlphabet command
It is often useful to be able to check if two alphabets are equivalent. The command:
1SOFIADecomposeAlphabet[alphabet1, alphabet2]
will check if (the differential of) each letter in the list alphabet1 can be decomposed linearly in terms of that of alphabet2. The input format for the first entry alphabet1 is a list of logarithms (for an explicit example, see knownAlphabet in Sec. 4.1). The second entry alphabet2 should have the same format as the output of the SOFIA[] command when FindLetters -> True, that is alphabet2 = { { Even letters} , { Odd letters} , { Square roots} } . Consequently, one can directly set $\mathtt{alphabet2=SOFIA}$[...]. This is useful to check whether the output of SOFIA[] spans an alphabet in the case that it is already known.
3.3.3 SOFIABaikov command
A loop-by-loop Baikov representation minimizing $N^{\prime}$ can be generated using:
1SOFIABaikov[{edges, nodes}, LoopEdges -> Automatic
2 , Dimension -> dim
3 , MaxCut -> False
4 , ShowXs -> False
5 , ShowDetailedDiagram -> True]
The options in lines 1 and 3 work the same way as the options with the same names in the SOFIA[] command. To set the spacetime dimension, one uses Dimension -> dim. When restricting to an integer dimension (e.g., $\mathtt{dim}=4$), then either (2.7) or its maximal cut (2.8) might vanish. This can happen because the prefactors include $\Gamma$-functions that may diverge; see, e.g., (A.21), which could indicate a potential $\frac{\varepsilon}{\varepsilon}$ cancellation. In such cases, [Uncaptioned image] will automatically display a warning and suggest to use dimensional regularization (e.g., $\mathtt{dim}=4-2\varepsilon$).
The Baikov variables, chosen internally by the code, can be printed in terms of loop and external momenta by setting ShowXs -> True. Enabling ShowDetailedDiagram -> True additionally displays these variables on the diagram specified by { edges, nodes} , along with the chosen internal edge momenta.
3.3.4 Graphical functions
[Uncaptioned image] offers several functions that leverage graph-theoretical techniques to streamline and accelerate singularity analysis, which may also be useful for other purposes. In particular:
1FeynmanDraw
2FeynmanPlot[{edges,nodes}, EdgeLabels -> {},
3 , HighlightedEdges -> {}]
4Subtopologies[{edges,nodes}, Reflections -> False]
5Symmetry[{edges1,nodes1},{edges2,nodes2}]
6SymmetryQuotient[ListOfDiagrams]
Drawing diagrams.
[…]
Plotting diagrams.
A tuple of the form { edges, nodes} can be verified to represent the correct Feynman diagram by running FeynmanPlot[{ edges, nodes} ]. The options EdgeLabels and HighlightedEdges allow labeling each element of edges (for instance, with edge momenta or Baikov variables) and highlighting specific edges by specifying their positions in edges. By default, if no options are provided, internal edges are decorated with the masses from edges and external edges with the masses from nodes.
Generating subtopologies.
Given a Feynman diagram represented as { edges, nodes} , the command Subtopologies[{ edges, nodes} ] generates a list of edges and nodes for each non-trivial subtopology of the original graph. It automatically removes contact diagrams, one-vertex reducible diagrams and tadpoles from the list of diagrams produced. For example, applying this command to any two-loop diagram will return a list of its one- and two-loop subtopologies, where the former arise from tadpole excisions—i.e., by removing internal edges of the form { { i,i} ,_ } (see for example Fig. 2(a)).
[…]
Symmetries.
Given two Feynman diagrams represented as { edges1, nodes1} and { edges2, nodes2} , the command Symmetry[{ edges1,nodes1} ,{ edges2,nodes2} ] identifies symmetry transformations relating them, accounting for masses and external momenta. If no symmetry exists, it returns an empty list: { } . Otherwise, it outputs { { edges,nodes} ,rule} where { edges, nodes} is one of the input diagrams and rule is the minimal replacement rule on the masses and Mandelstam variables required to relate it to the other diagram. In the above context, a reflection refers to a symmetry where rule = { } .
For example, Fig. 3 displays two pairs of diagrams (a-b and c-d) related by symmetry. Running Symmetry[(a),(b)]—with (a) and (b) representing the lists of edges and nodes of the diagrams of the same names—gives { (a),{ $\mathtt{s_{1,2}\!\to\!s_{2,3},s_{2,3}\!\to\!s_{1,2},M_{1}^{2}\!\to\!M_{2}^{2},M_{4}^{2}\!\to\!M_{1}^{2}}$} } , which corresponds to a clockwise rotation of $90^{\circ}$ relating diagram (a) to diagram (b).
Diagrams (c) and (d) have a less obvious relation. Running Symmetry[(c),(d)] outputs { (d),{ $\mathtt{s_{1,2}\!\to\!0,s_{2,3}\!\to\!0,M^{2}\to s_{1,2}}$} } which effectively amounts to setting the momentum $p_{4}$ of diagram (d) to zero, $p_{4}\to 0$. Indeed, in terms of Mandelstam variables, we have $\,p_{4}\cdot p_{1}=\,p_{2}\cdot p_{3}\propto s_{23}\to 0$ and $\,p_{4}\cdot p_{3}=\,p_{1}\cdot p_{2}\propto s_{12}\to 0$. In this case, Symmetry[] also correctly recognizes that the mass scale $M^{2}$ in diagram (d) plays the role of the Mandelstam invariant $s_{34}=s_{12}$ in diagram (c).
Figure 3: Diagrams related by the Symmetry function as described in the text. These diagrams were all generated by the FeynmanPlot function. The dashed lines denote massless particles and otherwise the mass label is displayed on the edges.
Note that the symmetry relation between diagrams is not necessarily bi-directional, as the example (d) $\to$ (c) demonstrates. We observe that Symmetry[] typically returns the diagram with the larger number of kinematic scales among the input diagrams. Furthermore, Symmetry[] can relate diagrams with different numbers of external edges. This includes cases where external momenta meet at a common vertex or where an external leg effectively “detaches” by setting its momentum exactly to zero.
In such cases, kinematic replacements can be subtle because the operations of specializing kinematics and computing singularities generally do not commute (a phenomenon analogous to that encountered when specializing kinematics and evaluating Feynman integrals; see [66, 67, 15]). To address this, we introduce a small parameter $\delta$ and extract the leading non-vanishing coefficient for each component of the singular locus $\mathcal{S}$ as $\delta\to 0$. See App. C for an explicit demonstration.
[…]
3.3.5 Solvers
All the documented functions so far have been specifically designed for applications to Feynman integrals. However, the solvers used behind the scenes can handle a variety of other problems, some of which are explored in Sec. 4. They can be used independently of the earlier commands via:
1SolvePolynomialSystem[polynomials, variables, Solver -> FastFubini
2 , SolverBound -> 100
3 , FactorResult -> True
4 , PLDMethod -> sym
5 , PLDHomogeneous -> true
6 , PLDHighPrecision -> false
7 , PLDCodimStart -> -1
8 , PLDFaceStart -> 1
9 , PLDRunASingleFace -> false]
Here, instead of the diagram information, the inputs consist of a set of polynomials and a set of variables to eliminate. The available options remain the same as above. In the case where Solver -> momentumPLD, the user must ensure that the variables to be eliminated are given as a text string in the form variables = { x1, x2, x3, ..., xn} . These variables are used to parametrize polynomials and are the default variables employed by the [Uncaptioned image] wrapper for Julia, i.e., they are predefined in the Julia external evaluator and not using them would cause failure. While the SOFIA function ensures this automatically when called, this step must be handled manually when using SolvePolynomialSystem.
4 Detailed examples
Figure 4: Example diagrams considered in Sec. 4. These diagrams were all generated by the FeynmanPlot function. The dashed lines denote massless particles and otherwise the mass label is displayed on the edges.
We now demonstrate the explicit use of the functions introduced above with concrete examples that are not too large to fit within the text. Collectively, these examples took less than 30 seconds to run on a conventional laptop. A list of 30 additional examples, including many multi-scale and multi-loop ones, can be found in the Mathematica notebook SOFIA_ examples.nb [3].
4.1 Symbol alphabet
In this example, we illustrate how [Uncaptioned image] , combined with the publicly available package Effortless [36], can be used to determine a (super)set of logarithmic letters, both even and odd, for a given Feynman integral.
We consider the diagram in Fig. 4(a), which contributes to the $t$-channel planar Møller and Bhabha scattering amplitudes at $\mathcal{O}(e^{6})$ in QED. The massive electrons and/or positrons are represented by solid lines, while the massless photon exchanges are depicted by dashed lines. This diagram was computed explicitly using differential equations in [68] and thus serves as an excellent crosscheck to the [Uncaptioned image] implementation.
A candidate alphabet can be generated by defining the graph in terms of its nodes and edges and then running the SOFIA[] command as follows:
1edges = {{{1,2},0}, {{2,5},m}, {{3,5},0}, {{3,4},m},
2 {{4,6},0}, {{1,6},m}, {{5,6},m}};
3nodes = {{1,m}, {2,m}, {3,m}, {4,m}};
4extraRoots = {-s12, -s23, -s12-s23};
5SOFIAalphabet = SOFIA[{edges, nodes}, FindLetters -> True
6 , SolverBound -> Infinity
7 , DoubleRoots -> {extraRoots, All}]
The output is a (super)set of 10 even and 42 odd candidate letters, which we collect in the alphabet SOFIAalphabet[[1;;2]]=$=\{\{\tilde{W}_{1},\ldots,\tilde{W}_{10}\},\{\tilde{W}_{11},\ldots,\tilde{W}_{52}\}\}$. We use the notation $\tilde{W}_{i}=\log(W_{i})$. For example:
$\tilde{W}_{31}$ = Log[$\frac{\mathtt{s12\,(4mm^{2}+mm\,s12-s12\,s23)+r[[5]]\,r[[10]]}}{\mathtt{s12\,(4mm^{2}+mm\,s12-s12\,s23)-r[[5]]\,r[[10]]}}$];
$\tilde{W}_{32}$ = Log[$\frac{\mathtt{s12\,(4mm-s12)(3mm-s23)+r[[5]]\,r[[10]]}}{\mathtt{s12\,(4mm-s12)(3mm-s23)-r[[5]]\,r[[10]]}}$];
3(* 50 more odd and even candidate letters *)
where the $\mathtt{r[[i]]}$’s are square roots, which will be discussed below. The $\tilde{W}_{i}$ depend on $\mathfrak{s}=3$ variables: $\mathtt{s12}$, $\mathtt{s23}$, and $\mathtt{mm}$. We introduce $\mathtt{mm}=\mathtt{m}^{2}$ here because SOFIA internally homogenizes the output polynomials and thus automatically replaces the mass variables $\mathtt{m}$ with the square root of their squares $\sqrt{\mathtt{mm}}$. In doing so, all the variables in the letters have the same mass dimension. In general, and as we will demonstrate explicitly below, not all of the letters in $\mathtt{SOFIAalphabet}$ are expected to appear in the differential equation (1.1).
It should also be noted that, in the example shown in Fig. 4(a), if the odd letters are assumed to be constructible only from the square roots of individual candidate singularities—as is the case with the default value DoubleRoots -> { } turned on—the resulting alphabet would be incomplete.
To be more explicit, let us clearly distinguish the effect of working with the option DoubleRoots -> { extraRoots,All} instead of the default one. Running:
1SOFIA[{edges, nodes}, (* same options as before *), DoubleRoots -> {}]
returns a set of odd letters that may depend on any of the 10 candidate square roots:
1r0s = {Sqrt[mm],
2 Sqrt[4 mm^3 - mm^2 s12 + 2 mm s12 s23 - s12 s23^2],
3 (* 8 more single square roots *)};
found by SOFIA, but where each odd letter contains at most one square root. Now, keeping the same settings but with the replacement DoubleRoots -> { extraRoots, All} , SOFIA will (internally) add to this initial list of 10 allowed roots $10\times 3=30$ double roots:
1r = Flatten[{Map[Sqrt[extraRoots] #&, r0s], r0s}];
for a total of 40 (single and double) square roots to use in the construction of $\mathbb{A}^{\text{odd}}$:
1r[[5]] = Sqrt[-s12] Sqrt[4mm-s12];
2r[[10]] = Sqrt[-s12] Sqrt[4 mm^3 - mm^2 s12 + 2 mm s12 s23 - s12 s23^2];
3(* 38 more single and double square roots *)
As we shall show momentarily, $\mathtt{r}$ is sufficient for the SOFIA command to produce a complete basis of odd letters. (Note that the signs chosen in the DoubleRoots option above are purely conventional and ensure that the comparison below is performed using the root conventions of [68]).
At this stage, we can check the validity of SOFIAalphabet by comparing it against the alphabet knownAlphabet in [68, Eq. (2.8)]. This is done by running:
1knownAlphabet = Log[$\frac{\mathtt{s12}}{\mathtt{mm}}$, $\frac{\mathtt{s12\,s23+r[[5]]\,r[[16]]}}{\mathtt{s12\,s23-r[[5]]\,r[[16]]}}$, (* 14 more letters *)}];
2SOFIADecomposeAlphabet[knownAlphabet, SOFIAalphabet]
In line 1 above, the syntax reflects that the roots defined in [68, Eq. (2.6)] were manually re-expressed, for comparison purposes, in terms of the roots $\mathtt{r[[i]]}$. The SOFIADecomposeAlphabet command then checks whether it is possible to decompose each element of knownAlphabet in terms of SOFIAalphabet and returns the decomposition if it is. For example, the most complicated letter in knownAlphabet decomposes according to:
$\mathrm{d}\log\frac{r_{5}r_{10}\left(m^{2}{-}s_{12}\right){-}4m^{6}s_{12}{-}3m^{4}s_{12}^{2}{+}m^{2}s_{12}^{3}{+}3m^{2}s_{12}^{2}s_{23}{-}s_{12}^{3}s_{23}}{r_{5}r_{10}\left(m^{2}{-}s_{12}\right){+}4m^{6}s_{12}{+}3m^{4}s_{12}^{2}{-}m^{2}s_{12}^{3}{-}3m^{2}s_{12}^{2}s_{23}{+}s_{12}^{3}s_{23}}=\mathrm{d}\tilde{W}_{31}{-}2\mathrm{d}\tilde{W}_{32}\,.$ (4.1)
See the Mathematica notebook SOFIA_ examples.nb [3] for all the details of the complete decomposition, as well as how the roots $r_{i}=\mathtt{r[[i]]}$ in r and the letters $\tilde{W}_{j}$ in SOFIAalphabet are defined. In the same notebook, we show that the same procedure successfully reproduces other non-trivial alphabets, including those arising in the two-loop massless and one-mass pentagon functions [69, 70, 71] as well as the genuine two-loop massless six-particle letters [72, 54].
4.2 Feynman geometries
In the following example, we examine the diagram shown in Fig. 4(b) with a top-quark loop of mass $m$. This diagram is particularly interesting because it contributes, through the double-box containing it, to the $\mathcal{O}(g_{s}^{4}m)$ amplitude for producing a Higgs boson of mass $M$ and a jet via gluon fusion at the LHC.
An important early step in computing this amplitude is to determine the types of functions to which the contributing diagrams are expected to evaluate, i.e., the function space. This information is encoded in the underlying geometry of the associated master integrals. To reveal it, it suffices to carefully analyze the maximal cut of each master integral in four dimensions ($\varepsilon=0$), which is conveniently done in the Baikov representation thanks to (2.8). This approach is reliable because the underlying geometry is characterized by the solution of the homogeneous part of the differential equation satisfied by the uncut integrals, which coincides with the differential equation governing the maximal cut integral [73, 74].
The maximal cut procedure always reduces Feynman integrals to periods over an algebraic variety. The geometry of this variety dictates the class of functions that appear after integration. For instance, a genus-0 (rational) curve gives rise to polylogarithms, a genus-1 (elliptic) curve to elliptic functions, and more intricate cases may involve higher-genus Riemann surfaces (see, e.g., [75]) or even higher-dimensional Calabi-–Yau manifolds (see Sec.,4.3), pointing to increasingly rich—and still not fully understood—mathematical structures. Strictly speaking, this way, one can only place an upper bound on the complexity of a given Feynman geometry; verifying that it is not degenerate generally requires further analysis. (To our knowledge, there is currently no automated algorithm to systematically determine the precise function space of a Feynman integral, e.g., whether it is polylogarithmic, elliptic, or beyond. Even less is known at the amplitude level, where full cancellations among non-polylogarithmic sectors may occur; see, e.g., [76]).
For the diagram in Fig. 4(b), we can use [Uncaptioned image] to learn about its geometry. Computing its maximal cut amounts to running:
1edges = {{{1,2},m}, {{2,3},m}, {{3,5},m}, {{4,5},m}, {{1,4},m}, {{2,5},0}};
2nodes = {{1,0}, {2,0}, {3,M}, {4,0}};
3SOFIABaikov[{edges, nodes}, MaxCut -> True, Dimension -> 4]
The result, obtained in a fraction of a second, is
[…]
The zero locus of the irreducible quartic in $x_{7}$, $P_{4}(x_{7})=\prod_{i=1}^{4}(x_{7}-r_{i})$, in the denominator defines the elliptic curve controlling the function space of this integral.
From (4.2), it is also easy to read off the leading singularities of the diagram in Fig. 4(b). Indeed, these are given by the zeroes of the discriminant of $P_{4}(x_{7})$ in $x_{7}$, i.e.,
$\begin{split}D_{x_{7}}P_{4}=m^{2}M^{2}s_{12}s_{23}&(M^{2}{-}s_{12})(M^{2}{-}s_{12}{-}s_{23})[16m^{4}(s_{12}{+}s_{23})^{2}{+}s_{23}^{2}(M^{2}{-}s_{12})^{2}\\ &{-}8m^{2}s_{23}(M^{2}(s_{23}{-}s_{12}){+}s_{12}(s_{12}{+}s_{23}))]\,.\end{split}$ (4.3)
These singularities can, of course, also be found automatically by running:
1SOFIA[{edges, nodes}, MaxCut -> True]
One can verify that all the singularities in the $\varepsilon$-factorized differential equation from [77] are included in the resulting list available in [3].
4.3 Eikonal/Regge/heavy-mass physics
This example considers the eikonal/Regge scattering of two heavy masses—e.g., black holes or neutron stars, treated as point particles—via graviton exchange. In particular, we focus on the diagram shown in Fig. 4(c), which contributes to the process at $\mathcal{O}(G_{N}^{5})$. This example demonstrates two key points: first, how to take the eikonal limit of the Baikov representation; and second, how the latter turns out sufficient to uncover the underlying Calabi–Yau threefold geometry [78, 42].
In order to take the eikonal/Regge/heavy-mass expansion and keep the leading order term which captures the relevant classical physics [79], we first parametrize the momenta in Fig. 4(c) such that
$\begin{split}p_{1}=\overline{p}_{1}-\frac{q}{2},\qquad p_{3}=-\overline{p}_{1}-\frac{q}{2},\qquad p_{2}=\overline{p}_{2}+\frac{q}{2},\qquad p_{4}=-\overline{p}_{2}+\frac{q}{2}\,.\end{split}$ (4.4)
Recall that we use all-incoming conventions for the momenta $p_{i}$. The spacelike momentum transfer $q$ satisfies the transverse condition $q\cdot\overline{p}_{i}\equiv 0$ for all $i$ as and is normalized so that $q^{2}\equiv-1$. Setting $\overline{p}_{i}=\overline{m}u_{i}$, with $u_{i}\cdot u_{i}\equiv 1$ and $u_{1}\cdot u_{2}\equiv y$, the loop-by-loop representation can be obtained in two steps.
First, we run the SOFIABaikov command on the diagram with the original kinematics. This can be done with:
1edges = {{{1,5},0}, {{2,5},0}, {{5,6},0}, {{5,7},0}, {{7,8},0}, {{3,7},0},
2 {{4,7},0}, {{1,6},m}, {{3,6},m}, {{2,8},m}, {{4,8},m}};
3nodes = {{1,m}, {2,m}, {3,m}, {4,m}};
4maxCut = SOFIABaikov[{edges, nodes}, MaxCut -> True
5 , Dimension -> 4
6 , ShowXs -> True]
The output maxCut is an expression for the maximal cut of this diagram. The option ShowXs prints also the definition of all the $x_{i}$’s, including the leftover integration variables $x_{12}$, $x_{13}$, $x_{14}$ we get after cutting all the $x_{1},x_{2},\ldots,x_{11}$. In particular, we find
$x_{12}=\ell_{4}^{2}\,,\qquad x_{13}=\ell_{3}^{2}\,,\qquad x_{14}=\ell_{4}\cdot p_{1}=\overline{m}\underbracket{\ell_{4}\cdot u_{1}}_{\hat{x}_{14}}+\ldots\,.$ (4.5)
The assignment of the loop momenta was automatic, but the precise knowledge of which momentum is which does not affect the analysis. We find:
$\mathtt{maxCut}=\frac{1}{2^{15}\pi^{17}}\int\frac{\mathrm{d}x_{12}\,\mathrm{d}x_{13}\,\mathrm{d}x_{14}}{\sqrt{x_{12}(4m^{2}-x_{12})}\sqrt{x_{13}(4m^{2}-x_{13})}\sqrt{P_{2}}\sqrt{P_{4}}}\,,$ (4.6)
where $P_{2}$ and $P_{4}$ are two polynomials with degrees $2$ and $4$ in the $x_{i}$’s respectively.
Implementing the eikonal limit is a matter of substituting the kinematic variables and expanding in large $\overline{m}=\mathtt{mb}$. It can be implemented with:
1eikonal = {s1,2 -> 2 (1 + y) mb^2,
2 s2,3 -> 2 (1 - y) mb^2,
3 m -> Sqrt[mb^2 - 1/4],
4 x14 -> x14h mb};
5maxCut = Series[Applyd[maxCut/.eikonal, {x12, x13, x14h}]
[…]
The list eikonal defines the substitutions, and the [Uncaptioned image] function Applyd implements the Jacobian for the change of variables to $(x_{12},x_{13},\hat{x}_{14})=(\mathtt{x12},\mathtt{x13},\mathtt{x14h})$. The leading order of this expansion gives:
$\mathtt{maxCut}=-\frac{1}{2^{19}\pi^{17}\overline{m}^{4}}\int\frac{\mathrm{d}x_{12}\,\mathrm{d}x_{13}\,\mathrm{d}\hat{x}_{14}}{\sqrt{x_{12}}\sqrt{x_{13}}\sqrt{\hat{P}_{2}}\sqrt{\hat{P}_{4}}}+\mathcal{O}(\overline{m}^{-5})\,,$ (4.7)
where the polynomials are given by
$\hat{P}_{2}$ $=4\hat{x}_{14}^{2}+(y^{2}-1)(1+x_{12})^{2}\,,$ (4.8)
$\hat{P}_{4}$ $=\hat{x}_{14}^{2}(1+x_{13})^{2}-x_{12}x_{13}(1+x_{12}+x_{13})\,.$ (4.9)
The curve $\hat{P}_{4}(x_{12},x_{13},\hat{x}_{14})=0$ defines a Calabi–Yau threefold (as a quartic polynomial in $\mathbb{CP}^{3}$ [80]), in agreement with the results of [78, 42].
The singularities of the resulting integral are simple to analyze. In [Uncaptioned image] , it boils down to running:
1SolvePolynomialSystem[{x12, x13, $\hat{P}_{2}$, $\hat{P}_{4}$}, {x12, x13, x14h}
2 , SolverBound -> Infinity]
The result gives singularities at $y=0$ and $y=\pm 1$, corresponding physically to the high-energy and static limits, respectively.
4.4 Beyond Feynman integrals
In this subsection, we illustrate how [Uncaptioned image] can be used to study singularities of other classes of Feynman-like integrals.
4.4.1 Energy-energy correlators
As an example, we focus on the type of integrals arising in the computation of the four-point energy-energy correlator (EEC) in the collinear limit within $\mathcal{N}=4$ super Yang-Mills theory [81]. Physically, the EEC captures meaningful correlations of energy flux after the emission of four final-state particles from an initial energetic parton and is now an active topic of research in both theoretical and experimental high-energy physics, see, e.g.[82, 83].
For four-point EEC, the relevant integrals to compute originate from a phase-space integral over a squared $1\to 4$ amplitude [81, Eqs. (2-4)]. For our discussion, it is sufficient to quote the two families of integrals considered in [81]:
$A_{a_{1}\dots a_{7}|a_{8}a_{9}a_{10}}$ $\equiv\int_{0}^{\infty}\mathrm{d}x_{1}\mathrm{d}x_{2}\mathrm{d}x_{3}\mathrm{d}x_{4}\frac{\delta[1-(x_{1}+x_{2}+x_{3}+x_{4})]}{s_{1234}^{a_{1}}s_{123}^{a_{2}}s_{234}^{a_{3}}x_{1234}^{a_{4}}x_{234}^{a_{5}}x_{123}^{a_{6}}x_{34}^{a_{7}}s_{12}^{a_{8}}s_{23}^{a_{9}}s_{34}^{a_{10}}}\,,$ (4.10a)
$B_{a_{1}\dots a_{4}|a_{5}a_{6}}$ $\equiv\int_{0}^{\infty}\mathrm{d}x_{1}\mathrm{d}x_{2}\mathrm{d}x_{3}\frac{\delta[1-(x_{1}+x_{2}+x_{3})]}{s_{123}^{a_{1}}x_{12}^{a_{2}}x_{23}^{a_{3}}x_{123}^{a_{4}}s_{12}^{a_{5}}s_{23}^{a_{6}}}\,,$ (4.10b)
where $a_{i}\in\mathbb{Z}$ and the variables entering (4.10) are
$s_{l_{1}l_{2}\cdots l_{n}}\equiv\sum_{1\leqslant i<j\leqslant n}x_{l_{i}}\,x_{l_{j}}\,(z_{l_{i}}-z_{l_{j}})\,(\bar{z}_{l_{i}}-\bar{z}_{l_{j}})\,,\qquad x_{l_{1}l_{2}\cdots l_{n}}\equiv\sum_{i=1}^{n}x_{l_{i}}\,,$ (4.11a)
$\text{with}\qquad(z_{1},\overline{z}_{1},z_{2},\overline{z}_{2},z_{3},\overline{z}_{3},z_{4},\overline{z}_{4})=\left(0,0,z,\overline{z},1,1,\frac{z-w}{1-w},\frac{\overline{z}-\overline{w}}{1-\overline{w}}\right)\,.$ (4.11b)
We treat $z$, $\bar{z}$, $w$, $\bar{w}$ as independent variables. The parametrization of these families is overcomplete since, e.g., $a_{2}$ and $a_{3}$ in (4.10) cannot be simultaneously positive. This would introduce an overlapping pole $\frac{1}{x_{12}x_{23}}$ which is forbidden by a version of Steinmann relations. We thus find it simpler to perform the singularity analysis at the level of master integrals introduced in [81, App. B] (in the same way we typically do the analysis at the level of individual Feynman diagram topologies, as opposed to all integrals at a given loop order at the same time).
The complete analysis is given in [3]. Below, we illustrate the procedure on a simple master integral
$B_{1}\equiv|z_{1}-z_{2}|^{2}\int_{0}^{\infty}\mathrm{d}x_{1}\mathrm{d}x_{2}\mathrm{d}x_{3}\frac{x_{1}\delta[1-(x_{1}+x_{2}+x_{3})]}{s_{123}\,x_{123}}\,.$ (4.12)
As explained in Sec. 2.2, to find the singular locus of this master integral, we can simply apply the Fubini reduction algorithm to the system of polynomials
=$\mathcal{L}=\{s_{123},x_{123},x_{1},x_{2},x_{3}\}\,,$ (4.13)
which explicitly can be inputted as:
1L = {z zb (x1 x2 + x2 x3) + x2 x3 (1 - z - zb) + x1 x3,
2 x1 + x2 + x3,
3 x1,
4 x2,
5 x3};
Note that $x_{1}\,,x_{2}$ and $x_{3}$ are included in $\mathcal{L}$ since $x_{1}=x_{2}=x_{3}=0$ appear in the integration boundaries of the $B$ family in (4.10b). Within [Uncaptioned image] , the singularities are computed by running:
1SolvePolynomialSystem[L, {x1, x2, x3}, SolverBound -> Infinity]
The output is
[…]
Looping this command over all the master integrals in [81, App. B] immediately reproduces the polynomial part of the alphabet quoted in that reference, namely
$\begin{split}\{z\,,&\overline{z}\,,1-z\,,1-\overline{z}\,,z-\overline{z}\,,w\overline{z}-\overline{w}z\,,w\overline{z}-\overline{w}\,,z\overline{w}-w\,,1-w-\overline{w}+z\overline{w}\,,\\ &1-w-\overline{w}+w\overline{z}\,,|z|^{2}-|w|^{2}\,,z-|w|^{2}\,,\overline{z}-|w|^{2}\}\cup(z\leftrightarrow w)\,.\end{split}$ (4.15)
Notably, the analysis over the master integrals
$\{A_{22},A_{23}\}\equiv|z_{2}-z_{3}|^{4}\int_{0}^{\infty}\mathrm{d}x_{1}\mathrm{d}x_{2}\mathrm{d}x_{3}\mathrm{d}x_{4}\frac{\{x_{2},x_{3}\}\delta[1-(x_{1}+x_{2}+x_{3}+x_{4})]}{s_{123}\,s_{234}\,x_{1234}}\,,$ (4.16)
returns a distinctly complicated letter (see SOFIA_ examples.nb [3] for details):
$L=w^{4}\overline{w}^{2}z^{2}+2w^{3}\overline{w}^{3}z^{2}-2w^{3}\overline{w}^{2}z^{2}+w^{2}\overline{w}^{4}z^{2}-2w^{2}\overline{w}^{3}z^{2}-2w^{2}\overline{w}^{2}z^{3}+(\text{151 terms})\,.$ (4.17)
It has an interesting geometric interpretation. As explained in [81], the solutions to the maximal cut conditions of the integrals in (4.16) (i.e., $s_{123}=s_{234}=x_{1234}=0$) define a singular cubic curve after projection onto the $(x_{2},x_{3})$-plane. Its roots $a\,,b\,,c$, which give rise to the non-integer polynomial part of the alphabet listed in [81, Eq. (12)], satisfy
$F=\Big{\{}abc=-|z|^{2}|w|^{2}\,,\,a+b+c=1-z-\overline{z}-w-\overline{w}\,,\,\frac{1}{a}+\frac{1}{b}+\frac{1}{c}=1-\frac{1}{z}-\frac{1}{\overline{z}}-\frac{1}{w}-\frac{1}{\overline{w}}\Big{\}}\,.$ (4.18)
The discriminant $D_{X}V$ of the cubic polynomial $V$ formed by these roots (given by Vièta’s formula)
$V=X^{3}-F_{2}X^{2}+F_{1}F_{3}X-F_{1}=0\,,$ (4.19)
is precisely (4.17):
$D_{X}V=[(a{-}b)(b{-}c)(c{-}a)]^{2}=L\,.$ (4.20)
4.4.2 Cosmological correlators
Similar singularity analysis can also be applied to cosmological correlators. Let us consider an example of a triangle diagram (also known as the one-loop three-site graph) for the conformally coupled scalar in Friedmann–Lemaître–Robertson–Walker cosmology. Following the notation of [84], the corresponding correlator expressed in Baikov representation is proportional to
$\int\frac{y_{12}y_{23}y_{31}\mathrm{d}y_{12}\,\mathrm{d}y_{23}\,\mathrm{d}y_{31}}{q_{\mathfrak{g}_{1}}q_{\mathfrak{g}_{2}}q_{\mathfrak{g}_{3}}}\kappa^{\chi}\left[\frac{1}{q_{\mathcal{G}_{12}}}\left(\frac{1}{q_{\mathfrak{g}_{23}}}+\frac{1}{q_{\mathfrak{g}_{31}}}\right)+\text{cyclic}\right]\,,$ (4.21)
where
$q_{\mathfrak{g}_{j}}$ $=x_{j}+y_{j-1,j}+y_{j,j+1}\,,$ (4.22a)
$q_{\mathfrak{g}_{j,j+1}}$ $=x_{j}+x_{j+1}+y_{j-1,j}+y_{j+1,j+2}\,,$ (4.22b)
$q_{\mathcal{G}_{j,j+1}}$ $=x_{1}+x_{2}+x_{3}+y_{j,j+1}\,,$ (4.22c)
with the indices defined modulo $3$. In addition, $\kappa$ is proportional to the Cayley–Menger determinant
$\kappa\propto\det\begin{bmatrix}0&1&1&1&1\\ 1&0&y_{12}^{2}&y_{23}^{2}&y_{31}^{2}\\ 1&y_{12}^{2}&0&x_{2}^{2}&x_{1}^{2}\\ 1&y_{23}^{2}&x_{2}^{2}&0&x_{3}^{2}\\ 1&y_{31}^{2}&x_{1}^{2}&x_{3}^{2}&0\end{bmatrix}\,,$ (4.23)
and $\chi$ is a generic parameter related to the type of cosmology. Physically, the $x_{i}$’s are variables measuring energies of the external particles.
The singularity analysis of this integral is straightforward. We obtain it by considering all $6$ terms in the square brackets of (4.21) separately. We find 37 possible singularities, for example:
$L=x_{1}^{4}+2x_{1}^{3}x_{2}-x_{1}^{2}x_{3}(2x_{2}+x_{3})-2x_{1}(x_{2}-x_{3})(x_{2}+x_{3})^{2}+2x_{3}(x_{2}+x_{3})^{3}\,,$ (4.24)
and record the rest in SOFIA_ examples.nb [3].
To our knowledge, the singularity set for the triangle integral was never obtained before. Partial results exist [85] and we have checked that our singularity set contains them.
5 Future directions
In this work, we introduced the package [Uncaptioned image] , whose main purpose is to automate the computations of singularities of Feynman integrals. It combines a new theoretical understanding of Baikov representations with advanced polynomial reduction techniques and graph-theoretical tools.
This first implementation already pushes beyond the current state of the art in many cases. In particular, the ancillary file [3] includes several new predictions, including those of Fig. 1. These results highlight both the strength and versatility of [Uncaptioned image] , leading to results that were previously beyond reach.
This progress also opens up a number of future directions in automatizing the determination of the analytic structure of scattering amplitudes and other objects of interest, such as cosmological or energy correlators.
Nature of singularities.
One of the most pressing questions is the algorithmic determination of the nature of singularities, i.e., the exponents $A$ and $B$ in the local behavior $f_{j}^{A}\log^{B}f_{j}$ as $f_{j}\to 0$. These exponents depend on the spacetime dimension and the geometry of the pinch manifold. In the case where the masses are sufficiently generic, this geometry is simple to describe, and consequently a closed-form formula for the exponents $A$ and $B$ is known [9]; see, e.g., [44, 86] for reviews. In “non-generic” situations, which are often the ones of physical interest, more work has to be done in analyzing the pinch manifold. Performing this analysis algorithmically would provide a significant improvement over the capabilities of [Uncaptioned image] , especially given that it would allow one to distinguish between square-root and logarithmic letters, thus optimizing the size of the proposed symbol alphabet.
Sheet structure.
A related question concerns the Riemann sheet structure of Feynman integrals and scattering amplitudes. The simplest version of this question is to determine which singularities lie on the physical sheet or in the physical region. This question is usually addressed by either determining whether the loop momenta along the pinch manifold are physical, or whether the associated Schwinger parameters are non-negative. We have not attempted to address these questions since the Fubini reduction algorithm circumvents the need to solve the Landau equations directly. Therefore, it does not directly reconstruct the pinch manifold. Similarly, the discontinuity (monodromy) structure of Feynman integrals is more difficult to study in Baikov-like representations, as has been recognized for a long time [38, 87], since the integration contour itself depends on the kinematics. One possible direction is to keep track of the genealogy of singularities in the reduction algorithm, see [32, Sec. 6], which is closely related to the hierarchical structure; see, e.g., [88].
The questions about the sheet structure have their avatar in the symbol properties for polylogarithmic integrals. For example, presence on the physical sheet corresponds to the first-entry condition and the discontinuity structure is encoded in symbol adjacency. See [89, 90] for recent work on the connection to Landau analysis. We believe that the same information is encoded in the commutation relations of the matrices $\mathbf{\Omega}_{i}$ in (1.1) and boundary conditions even for non-polylogarithmic integrals, which means being able to predict it independently would put constraints on the $\mathbf{\Omega}_{i}$’s. See [70, 91] for explicit examples in the context of extended Steinmann relations.
Finding the alphabet.
To find the symbol alphabet we wrapped in [Uncaptioned image] the publicly available package Effortless [36]. More work is needed in the algorithmic determination of alphabets, which is currently the main bottleneck in [Uncaptioned image] as can be seen in the timings indicated in Fig. 1. We have already mentioned that this procedure would benefit from determining algorithmically which letters are square-root vs. logarithmic, though this information can be provided through the option AddLetters, as explained in Sec. 3.3.1. In addition, more conceptual work is needed to understand how general letters can be, including the maximum multiplicity of square roots that can be expected and which singularities are allowed to appear as root arguments. Here, we committed to the ansatz (2.36) with double square roots. More general ansätze can be implemented as needed.
Feynman geometries.
We found the optimized loop-by-loop Baikov representation to be efficient at bounding the complexity of the geometries associated with Feynman integrals, e.g., whether they are expressible in terms of periods of elliptic curves or Calabi–Yau manifolds. In addition, [Uncaptioned image] allows one to find the loci of parameters for which they become singular. One may imagine extending [Uncaptioned image] to also analyze these geometries completely automatically.
Spacetime dimension.
In the current implementation, [Uncaptioned image] assumes that the spacetime dimension $\mathrm{D}$ is large enough so that all external momenta remain independent up to momentum conservation. It also assumes that $\mathrm{D}$ is generic, which is what is needed for dimensional regularization. This is why we only provide candidate singularities. Whenever one specializes the dimension $\mathrm{D}$, the actual list of singularities might, and in general will, reduce. Therefore, another improvement on [Uncaptioned image] would be to analyze such cases separately. Imposing a constraint on the external momenta is cumbersome because one has to work with constrained variables, or come up with a different set of kinematical data such as spinors. On the other hand, imposing a bound on the dimensionality of the loop momenta could potentially reduce the number of integration variables in Baikov representations.
Specializing the spacetime dimension may prove useful in other contexts, such as the S-matrix bootstrap [92], where understanding the singularity structure can be essential to make use of tools such as dispersion relations, particularly at higher points [13, 93] and beyond lightest-particle scattering [94, 95]. Another relevant application is the study of two-dimensional integrable models, in which anomalous thresholds manifest as Coleman–Thun poles [96, 97]. For example, an automated tool like [Uncaptioned image] could help extend recent studies on perturbative integrability beyond one-loop [98, 99].
Bottleneck diagrams.
We observe that the performance of [Uncaptioned image] could be further optimized for Feynman diagrams that exhibit maximal connectivity—that is, diagrams for which the minimum number of vertex removals required to disconnect the graph into trees is maximized. This pattern becomes evident when monitoring the SOFIA[] command on sufficiently complicated diagrams with IncludeSubtopologies -> True: the (sub)topologies taking the highest runtime are those that are maximally, or nearly maximally, connected.
One main drawback of the loop-by-loop Baikov approach in these cases is that the resulting $N$ in (2.6) typically remains high, because each $\mathcal{E}_{a}$ is as large as possible for maximally connected diagrams, regardless of our minimizing strategy for the choice of LoopEdges. This poses a computational challenge starting at three-loops, because maximally connected diagrams lead to systems with a large number ($\geqslant 5$) of variables to eliminate, even on the maximal cut. An example (with $N=5$ on the maximal cut) is the symmetric “envelope diagram” from [66, Fig. 1e], whose singularities were previously calculated with PLD.jl in Schwinger parameter space for different masses assignments. Although the symbolic extraction of singularities for this diagram currently requires an impractical runtime in [Uncaptioned image] , we were able to use it to successfully reproduce numerically all of its known singularities.
We emphasize that these diagrams remain equally challenging to solve with either Solver -> FastFubini or Solver -> momentumPLD in SOFIA[]. For the envelope, the difficulty in the latter arises because, on the max cut, one must still eliminate five Baikov variables (the $x$’s) and five $\beta$’s from (2.23) (with one $\beta$ fixed to 1). In total, this requires eliminating nine variables in Baikov space, compared to only five in Schwinger space.
Fortunately, the stringent criteria of maximal vertex connectivity ensure that, at a given loop order and multiplicity, the removal of only a few vertices results in most diagrams disconnecting into trees. Consequently, the number of maximally connected diagrams remains combinatorially small, especially for processes of phenomenological relevance, e.g., in the Standard Model.
Finally, because the connectivity requirement naturally forces a uniformity in the distribution of vertices, maximally connected diagrams are generally highly symmetric. Fully exploiting this property, as demonstrated in [34, 35] for the physical sheet singularity, could plausibly lead to further improvements of [Uncaptioned image] ’s performance on these graphs.
Acknowledgments.
We thank Samuel Abreu, Giacomo Brunello, Simon Caron-Huot, Hjalte Frellesvig, Hofie Hannesdottir, Antonela Matijašić, Andrew McLeod, Ian Moult, Erik Panzer, Andrzej Pokraka, Simon Telen, and Simone Zoia for useful comments and discussions. In particular, we thank Antonela Matijašić and Julian Miczajka for sharing pre-release versions of Effortless with us. The work of M.C. and M.G. is supported by the National Science and Engineering Council of Canada (NSERC) and the Canada Research Chair program, reference number CRC-2022-00421. M.G. thanks the Institute for Advanced Study for its hospitality during a crucial stage of this work.
Appendix A Derivation of Baikov representations
Setup.
We begin by expressing a generic Feynman integral $I$ in a form that emphasizes the relevant structure for the derivation, particularly the integration measure:
$I=\int\mathrm{d}^{\mathrm{D}}\ell_{1}\,\mathrm{d}^{\mathrm{D}}\ell_{2}\cdots\mathrm{d}^{\mathrm{D}}\ell_{L}\,\mathcal{I}(\ell_{1},\ell_{2},\dots,\ell_{L})\,,$ (A.1)
where the integrand $\mathcal{I}$ is a rational function of the loop and external momenta, $\ell_{a}$ and $p_{i}$ respectively, as well as possible other variables such as masses.
Let us use the label $q_{i}$ to collectively refer to the independent internal and external momenta, that is
$(q_{1},q_{2},\ldots,q_{M})=(p_{1},p_{2},\ldots,p_{\mathcal{E}},\ell_{L},\ldots,\ell_{2},\ell_{1})\,.$ (A.2)
We only include $\mathcal{E}$ external momenta, where
$\mathcal{E}=\min(\mathrm{D}_{0},n-1)\,.$ (A.3)
This is because there are at most $n-1$ independent $p_{i}$’s due to momentum conservation. In addition, the number of independent $p_{i}$’s cannot exceed the spacetime dimension $\mathrm{D}_{0}$ (in dimensional regularization, $\mathrm{D}=\mathrm{D}_{0}-2\varepsilon$). The total number of such independent momenta $q_{i}$ is therefore $M=L+\mathcal{E}$. They define a natural coordinate system, where we work in terms of the Lorentz-invariant scalar products: $X_{ij}=q_{i}\cdot q_{j}$.
Let us consider the integration measure $\mathrm{d}^{\mathrm{D}}\ell_{a}$ in the order $a=1,2,\ldots,L$. We can assume that all the momenta $\ell_{1},\ell_{2},\ldots,\ell_{a-1}$ have been handled and now we consider $\ell_{a}$. The first key observation is that $\ell_{a}$ can be decomposed into two components:
$\ell_{a}=\ell_{a\parallel}+\ell_{a\perp}\,\qquad(1\leqslant a\leqslant L)\,,$ (A.4)
where $\ell_{a\parallel}$ lies in the space spanned by $(q_{1},\dots,q_{M-a})=(p_{1},\ldots,p_{\mathcal{E}},\ell_{L},\ldots,\ell_{a+1})$, and $\ell_{a\perp}$ is perpendicular to this space. The integration measure then factorizes as
$\mathrm{d}^{\mathrm{D}}\ell_{a}=\mathrm{d}^{M-a}\ell_{a\parallel}\,\mathrm{d}^{\mathrm{D}-M+a}\ell_{a\perp}\,\qquad(1\leqslant a\leqslant L)\,.$ (A.5)
Our next goal is to change the variables in the parallel components $\ell_{a\parallel}$ of the loop momenta to the scalar products $X_{ij}$. In order to do so, we first derive the following general result.
Key equation.
Consider any vector $v$ in an $n$-dimensional space and express it in terms of a (not necessarily orthonormal) basis $(q_{1},q_{2},\dots,q_{n})$:
$v=\lambda_{1}\,q_{1}+\lambda_{2}\,q_{2}+\cdots+\lambda_{n}\,q_{n}\,.$ (A.6)
In components, this reads
$v^{i}=\sum_{a=1}^{n}\lambda_{a}\,q_{a}^{i}\,.$ (A.7)
The differential volume element in the $n$-dimensional subspace is then given by
$\mathrm{d}^{n}v=\Bigl{|}\det\!\frac{\partial v^{i}}{\partial\lambda_{a}}\Bigr{|}\,\mathrm{d}^{n}\lambda\,.$ (A.8)
Since $\frac{\partial v^{i}}{\partial\lambda_{a}}=q_{a}^{i}$, the corresponding Jacobian matrix $J$ has entries $J_{ia}=q_{a}^{i}$. Although the basis $(q_{1},...,q_{n})$ is not assumed orthonormal, we can compute
$(J^{T}J)_{ab}=q_{a}\cdot q_{b}\equiv G(q_{1},q_{2},\dots,q_{n})_{ab}\,,$ (A.9)
where $G=G(q_{1},q_{2},\dots,q_{n})$ is called the Gram matrix. Noting that $\det(J^{T}J)=(\det J)^{2}$, we deduce $(\det J)^{2}=\det G$ and hence: $\det J=\pm\sqrt{\det G}$. Thus, the change of variables yields
$\mathrm{d}^{n}v=\sqrt{\det G}\;\mathrm{d}^{n}\lambda\,.$ (A.10)
Note that the factor $\sqrt{\det G}$ is nothing but the volume of the parallelepiped spanned by the vectors $q_{1},\dots,q_{n}$. This factor accounts for the distortion of the unit hypercube in the $\lambda$-space when mapped into the momentum space by the basis $\{q_{a}\}$. For an orthonormal basis, $\det G=1$, and the volume is unchanged. For a non-orthonormal basis, the hypercube is transformed into a parallelepiped with volume $\sqrt{\det G}$, which corrects the integration measure accordingly. Equipped with (A.10), we proceed with the derivation.
Parallel components. The parallel components are, by definition, linear combinations of the basis vectors $q_{1},q_{2},\dots,q_{M-a}$:
$\ell_{a\parallel}=\sum_{b=1}^{M-a}\lambda_{ab}q_{b}\,\qquad(1\leqslant a\leqslant L)\,.$ (A.11)
Under the projection onto this $q$-basis, the volume form transforms according to (A.10):
$\begin{split}\mathrm{d}^{M-a}\ell_{a\parallel}&=\sqrt{\det G}\prod_{b=1}^{M-a}\mathrm{d}\lambda_{ab}\,,\end{split}$ (A.12)
where now $G=G(q_{1},q_{2},\dots,q_{M-a})$. Next, using (A.11), it is easy to see that the $X$’s and the $\lambda$’s are linearly related
$X_{ak}=q_{k}\cdot\ell_{a\parallel}=\sum_{b=1}^{M-a}\lambda_{ab}q_{b}\cdot q_{k}\iff X=\lambda\cdot G\,,$ (A.13)
where the right-hand side is a matrix equation. Thus,
$\begin{split}\mathrm{d}^{M-a}\ell_{a\parallel}&=\sqrt{\det G}\Bigl{|}\det\!\frac{\partial X}{\partial\lambda}\Bigr{|}^{-1}\prod_{b=1}^{M-a}\mathrm{d}X_{ab}=\frac{\prod_{b=1}^{M-a}\mathrm{d}X_{ab}}{\sqrt{\det G}}\,\qquad(1\leqslant a\leqslant L)\,.\end{split}$ (A.14)
Perpendicular components. The perpendicular components $\ell_{a\perp}$ are integrated in spherical coordinates, where the volume element in an $(\mathrm{D}-M+a)$-dimensional space is
$\mathrm{d}^{\mathrm{D}-M+a}\ell_{a\perp}=\ell_{a\perp}^{\mathrm{D}-M+a-1}\mathrm{d}\ell_{a\perp}\mathrm{d}\Omega_{\mathrm{D}-M+a}\,.$ (A.15)
By integrating out the angular part, we obtain
$\int\limits_{S_{\mathrm{D}-M+a}}\mathrm{d}^{\mathrm{D}-M+a}\ell_{a\perp}=\frac{1}{2}\Omega_{\mathrm{D}-M+a}\,\ell_{a\perp}^{\mathrm{D}-M+a-2}\,\mathrm{d}\ell_{a\perp}^{2}\,,$ (A.16)
where
$\Omega_{n}=\frac{2\pi^{n/2}}{\Gamma(n/2)}\,,$ (A.17)
is the full $n$-dimensional solid angle. Choosing the $(q_{1},\dots,q_{M-a})$ such that $\ell_{a\perp}^{2}=X_{aa}$, we substitute $\mathrm{d}\ell_{a\perp}^{2}=\mathrm{d}X_{aa}$, leading to
$\int\limits_{S_{\mathrm{D}-M+a}}\mathrm{d}^{\mathrm{D}-M+a}\ell_{a\perp}=\left[\frac{\det G(\ell_{a},\ell_{a+1},\dots,p_{\mathcal{E}})}{\det G(\ell_{a+1},\dots,p_{\mathcal{E}})}\right]^{\frac{\mathrm{D}-M+a-2}{2}}\frac{\Omega_{\mathrm{D}-M+a}\,\mathrm{d}X_{aa}}{2}\,,$ (A.18)
for $1\leqslant a\leqslant L$. This formula has an insightful geometric interpretation. The loop momentum $\ell_{a}$ is decomposed into two parts: one lying within the subspace spanned by $(\ell_{a+1},\dots,\ell_{L},p_{1},\dots,p_{\mathcal{E}})$ (base space) and the other orthogonal to it (ruled hypersurface). The Cartesian product of these two subspaces forms a parallelepiped. Just as the height of a three-dimensional cylinder can be described as the ratio of two volumes, the height/radius of the ruled hypersurface can similarly be expressed in terms of the volume ratio:
$|\ell_{a\bot}|\equiv\sqrt{X_{aa}}=\sqrt{\frac{\det G(\ell_{a},\ell_{a+1},\dots,\ell_{L},p_{1},\dots,p_{\mathcal{E}})}{\det G(\ell_{a+1},\dots,\ell_{L},p_{1},\dots,p_{\mathcal{E}})}}\,.$ (A.19)
As in (A.10), the Gram determinant $\det G(q_{1},\dots,q_{\mathrm{D}-M+a})$ gives the square of the volume of the parallelepiped spanned by the vectors $(q_{i})_{i=1}^{\mathrm{D}-M+a}$. Hence, the numerator in (A.19) represents the volume of the entire parallelepiped formed by $(\ell_{a},\ell_{a+1},\dots,\ell_{L},p_{1},\dots,p_{\mathcal{E}})$ while the denominator represents the volume of the base spanned by $(\ell_{a+1},\dots,\ell_{L},p_{1},\dots,p_{\mathcal{E}})$.
Finally, since the integration region for the radial coordinate $\mathrm{d}\ell_{a\perp}^{2}$ in (A.16) is $\ell_{a\perp}^{2}>0$, the integration region for the scalar measure in (A.18) is simply given by
$\Gamma_{a}\equiv\left\{\ell_{a\perp}^{2}=\frac{\det G(\ell_{a},\dots,\ell_{L},p_{1},\dots,p_{\mathcal{E}})}{\det G(\ell_{a+1},\dots,\ell_{L},p_{1},\dots,p_{\mathcal{E}})}>0\right\}\,\qquad(1\leqslant a\leqslant L)\,.$ (A.20)
Final expression. After substituting the parallel (see (A.14)) and perpendicular (see (A.18)) volume elements for each loop momentum, all but the $a=1$ and $a=L$ Gram determinants cancel, resulting in the Baikov representation of (A.1)
$\begin{split}\int\mathrm{d}^{\mathrm{D}}\ell_{1}\mathrm{d}^{\mathrm{D}}\ell_{2}\dots\mathrm{d}^{\mathrm{D}}&\ell_{L}\,\mathcal{I}(\ell_{1},\dots,\ell_{L})=\frac{\pi^{\frac{L(1+2(\mathrm{D}+\mathcal{E})-L)}{4}}}{\prod_{l=1}^{L}\Gamma\left(\frac{\mathrm{D}-M+l}{2}\right)}\det G(p_{1},\dots,p_{\mathcal{E}})^{(-\mathrm{D}+\mathcal{E}+1)/2}\\ &\times\int_{\Gamma}\prod_{i=1}^{L}\prod_{j=i}^{M}\mathrm{d}X_{ij}\det G(\ell_{1},\dots,\ell_{L},p_{1},\dots,p_{\mathcal{E}})^{(\mathrm{D}-M-1)/2}f(X_{ij})\,,\end{split}$ (A.21)
where the multiple factors of $i\pi^{\mathrm{D}/2}$ in (2.1) were absorbed in the integrand $f$. Moreover, the integration domain $\Gamma$ is given by imposing $L$ conditions given in (A.20), such that $\Gamma\equiv\Gamma_{1}\cap\Gamma_{2}\cap\cdots\cap\Gamma_{L}=\{\det G(\ell_{1},\dots,\ell_{L},p_{1},\dots,p_{\mathcal{E}})>0\}$. [^2] The last equality follows from the observation that, since $\det G(q_{1},q_{2},\dots,q_{n})>0$, the biggest Gram matrix is positive definite. A key property of positive definite matrices is that every principal submatrix is also positive definite. In particular, Sylvester’s criterion tells us that all leading principal minors are positive, and in fact, this extends to any principal submatrix. Therefore, the submatrix corresponding to $q_{2},q_{3},\dots,q_{n}$ is positive definite, ensuring that $\det G(q_{2},q_{3},\dots,q_{n})>0$ and so on. In contrast, in the loop-by-loop case discussed in the main text, the multiple pairs of internal-external Gram matrices typically are not principal submatrices of one another. This occurs when they satisfy (A.21), but with Gram matrices of different sizes (see, e.g., (2.15)). In such cases, the integration contour remains the complete intersection of the corresponding $\Gamma_{i}$ regions.
Note that the number of integration variables $X_{ij}$ in the measure product corresponds to the number of unique pairs $(i,j)$, where $i$ runs from $1$ to $L$ and $j$ runs from $i$ to $M$. This count evaluates to
$\sum_{i=1}^{L}(M-i+1)=\frac{L(L+1)}{2}+L\mathcal{E}\equiv N^{\prime}\,,$ (A.22)
when $M=L+\mathcal{E}$. In the main text, we relabeled the $N^{\prime}$ integration variables as
$(x_{1},\dots,x_{N^{\prime}})=\text{linear combinations of }\bigcup_{i=1}^{L}\{X_{ij}\mid i\leqslant j\leqslant M\}\,,$ (A.23)
where the first $E$ $x$’s are propagators. Given this notation, let us finally note that, in Sec. 2.1, we denoted
$c$ $=\frac{2^{L-N}(-i)^{L}\pi^{(L-N)/2}}{\prod_{l=1}^{L}\Gamma\left(\frac{\mathrm{D}-M+l}{2}\right)}\,,$ (A.24a)
$\gamma$ $=(\mathrm{D}-M-1)/2\,,$ (A.24b)
$\tilde{\gamma}$ $=(\mathcal{E}-\mathrm{D}+1)/2\,,$ (A.24c)
$\mathcal{G}(\vec{x};\vec{s})$ $=\det G(\ell_{1},\dots,\ell_{L},p_{1},\dots,p_{\mathcal{E}})\,,$ (A.24d)
$\tilde{\mathcal{G}}(\vec{s})$ $=\det G(p_{1},\dots,p_{\mathcal{E}})\,.$ (A.24e)
This completes the derivation of the global Baikov representation in (2.4).
Loop-by-loop variants. In contrast, the loop-by-loop Baikov representations (2.7) are derived from the global Baikov formula (A.21) by first choosing an ordering for the loops in the diagram considered. For each loop, one identifies the set of linearly independent external momenta (these include momenta from loops to be integrated later). Then, the one-loop Baikov representation (i.e., (A.21) for $L=1$) is applied to that loop momentum. Repeating this process for every remaining loop yields a representation that generally involves fewer integration variables than the global Baikov representation (see, e.g., the sunrise example discussed in Sec. 2.1). Note that this procedure is not unique—the chosen loop ordering and the selection of independent momenta can influence both the final number of integration variables $N$ and the form of the integrand. The package [Uncaptioned image] automatically looks for a choice of loop shifts and ordering that minimizes $N$.
Appendix B Interface with Julia
As mentioned in Sec. 3, [Uncaptioned image] can optionally use the Julia package PLD.jl to extract the singular locus in (2.24). A working installation of this package can be linked to [Uncaptioned image] with the option
1SOFIAoptionJulia = True;
Running this command successfully for the first time requires a few steps. First, the user must ensure that PLD.jl and all its dependencies are correctly installed and running on their Julia interface. The installation procedure can be found in [65].
Next, the user needs to register the version of Julia in which PLD.jl runs (i.e., v1.8.5) as an external evaluator in Mathematica by running
1RegisterExternalEvaluator["Julia", "/path/to/julia"]
where “/path/to/julia” is replaced by the actual path to the Julia binary. The user can then verify that Mathematica is communicating with Julia correctly by running, e.g., ExternalEvaluate["Julia", "1 + 1"]. Only once this is done, one should be able to run Get["SOFIA.m"] with SOFIAoptionJulia = True; if the script is loaded successfully, a “Checking status” log should appear while loading without warnings or errors on the Mathematica interface.
Appendix C Subtleties with kinematic replacements
As described in Sec. 3, in many cases the command Symmetry[] relates two diagrams with different numbers of scales. In such situations, some singularities might be missed if the kinematic replacement is not taken carefully.
To illustrate this point, we consider the two acnode diagrams defined by:
1(* Acnode diagram #1 *)
2edges1 = {{{1,2},m}, {{2,3},m}, {{3,4},0}, {{2,4},0}, {{1,3},m}};
3nodes1 = $\mathtt{M_{1}}$}, {2,$\mathtt{M}$}, {3,$\mathtt{M}$}, {4,$\mathtt{M}$}};
4(* Acnode diagram #2 *)
5edges2 = {{{1,2},0}, {{2,3},0}, {{3,4},0}, {{2,4},0}, {{1,3},0}};
6nodes2 = {{1,0}, {2,$\mathtt{M}$}, {3,$\mathtt{M}$}, {4,$\mathtt{M}$}};
We can plot them via:
1FeynmanPlot/@{{edges1,nodes1},{edges2,nodes2}}
which outputs the diagrams in Fig. 5, with diagram (a) corresponding to { edges1,nodes1} , and diagram (b) corresponding to { edges2,nodes2} .
To find the symmetry relation we run:
1Symmetry[{edges1,nodes1},{edges2,nodes2}]
confirming that diagram { edges1,nodes1} is related to { edges2,nodes2} via the kinematic replacement $r_{0}=${ $\mathtt{M_{1}^{2}}\to\mathtt{0},\mathtt{m^{2}}\to\mathtt{0}$} .
Figure 5: Diagrams related by the command Symmetry[] and whose kinematic replacement rule must be taken with care, as described in the text. These diagrams were generated by the FeynmanPlot[] function.
Let us now see how this replacement rule applied naively is insufficient to reproduce all singularities. It is enough to perform the leading singularity analysis with IncludeSubtopologies -> False on both diagrams:
1diags = {{edges1,nodes1},{edges2,nodes2}};
2Do[singularitiesAcn[i] = SOFIA[diags[[i]], SolverBound -> 600
3 , IncludeSubtopologies -> False]
4 , {i,2}];
Applying the naive replacement rule $r_{0}$ directly to the eleven components of $\mathcal{S}$ obtained in singularitiesAcn[1] yields all eight components of $\mathcal{S}$ found for singularitiesAcn[2], except for one:
$\mathtt{MM}^{2}-3\,\mathtt{MM}\,\mathtt{s12}+\mathtt{s12}^{2}+\mathtt{s12}\,\mathtt{s23}\,,$ (C.1)
where, we recall, $\mathtt{MM}=\mathtt{M}^{2}$, $\mathtt{MM1}=\mathtt{M}_{1}^{2}$ and $\mathtt{sij}=\mathtt{s}_{ij}$.
However, by replacing $r_{0}$ with its “blown-up” version, $r_{\delta}=$ $\mathtt{M_{1}^{2}}\to\mathtt{\delta},\mathtt{m^{2}}\to\mathtt{\delta}$} , expanding to leading order in $\mathtt{\delta}$, and factoring the result into irreducible components, we recover all the singularities in singularitiesAcn[2], including (C.1). In particular, from the $11^{\text{th}}$ component of $\mathcal{S}$ obtained for singularitiesAcn[1], we have
[…]
For all the details of this example, see the notebook SOFIA_ examples.nb in [3]. The $\delta$-procedure outlined above is implemented and executed automatically by [Uncaptioned image] .